### v5

# Algebroid functions on complex manifolds

Abstract.
We build a basic theory for algebroid functions on complex manifolds including several closely related topics such as uniformization problems, Nevanlinna theory, propagation of algebraic dependence and unicity problems, which is a development of the classical theory for algebroid functions on the complex plane. To the best of our knowledge, it is the first result for algebroid functions on a nontrivial complex manifold after H. Poincaré proposed the notion of algebroid functions.
Key words and phrases:
Algebroid functions; uniformization problems; second main theorem; Picard’s theorem; unicity.2010 Mathematics Subject Classification:
32H30; 32H25Contents
- 1 Introduction
- 2 Algebroid functions on Hermitian manifolds
- 3 Uniformization of algebroid functions on Hermitian manifolds
- 4 Nevanlinna theory of meromorphic functions on Kähler manifolds
- 5 Nevanlinna theory of algebroid functions on Kähler manifolds
- 6 Unicity problems of algebroid functions on Kähler manifolds
1. Introduction
1.1. Background
The notion of algebroid functions (see [22, 41]) introduced by H. Poincaré is a generalization of algebraic functions and meromorphic functions. Let us review that a -valued algebroid function on a (nonempty) domain (the complex plane) is defined by an irreducible algebraic equation
where are holomorphic functions with no common zeros on An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G. Darboux deemed that it is an important class of functions. To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a (generic) -sheeted analytic covering of A further fact claims that can lift to a single-valued meromorphic function on via a natural projection satisfying that it therefore realizes the uniformization of In fact, it leads to the famous uniformization problem for Riemann surfaces, which had been solved by Poincaré-Klein-Koebe (see [28, 35]) who showed the uniformization theorem which asserts that every Riemann surface is conformally equivalent to some quotient space where is one of the three canonical domains: or (the unit disc), while is a proper discrete group of Möbius automorphisms of
Even so, the most concerned problem on algebroid functions is their value distribution. About 1930s after R. Nevanlinna [33] founded two fundamental theorems concerning the value distribution theory of meromorphic functions in 1925, many famous scholars such as G. Rémoundos, H. Selberg, E. Ullrich and G. Valiron, etc. began to study the value distribution theory of algebroid functions. For example, G. Rémoundos [36] proved a Picard’s theorem which asserts a nonconstant -valued algebroid function on can omit distinct values at most in Afterwards, G. Valiron [59] showed without details a five-value type theorem in unicity theory which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to this topic, involving investigations to Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc.. More details refer to H. Cartan [9], He-Xiao [22], Niino-Ozawa [29], J. Noguchi [30], M. Ozawa [34], M. Ru [37, 38], Sun-Gao [41], J. Suzuki [43], L. Selberg [45, 46], N. Toda [54, 55], E. Ullrich [58] and their references therein.
The algebroid function theory in one complex variable has received a great deal of achievements, however, it didn’t mature until the last 30-40 years. So far, there is very little literature discussing the algebroid functions in several complex variables, except part of contributions of Hu-Yang [21] to the theory on due to some technical difficulties such as the complexity of branching points and multiple points as well as uniformization of an algebroid function. We even know nothing about the algebroid functions on a nontrivial complex manifold. Given two nonconstant algebroid functions on a non-compact Hermitian manifold (referred to (1) in Section 2.1 for definition). We are inspired by the following closely related fundamental problems:
How to realize the uniformization of ?
How many Picard’s exceptional values can have at most?
How many values do have to share at least to guarantee ?
To solve the first problem, we construct a multi-leaf complex manifold via which is a generic multi-sheeted analytic covering of Then, we lift to a meromorphic function on via the natural projection with The strategy of solving the second problem is to establish the Nevanlinna theory of algebroid functions on a complete Kähler manifold under certain curvature condition. In order to establish the Second Main Theorem of our trick used is to lift to via Thereout, one can receive the defect relation of which derives the Picard’s theorem. The central idea of dealing with the last problem is the applications of Nevanlinna theory. In doing so, we generalize for the first time the propagation theory of algebraic dependence introduced by L. Smiley [47] to the algebroid functions. This is a technical tool to study unicity problems. Based on it, we obtain the sharp lower bound of the number of the shared values of guaranteeing
1.2. Main results
Our study of algebroid function theory covers four closely related aspects: uniformization problems, Nevanlinna theory, propagation theory of algebraic dependence and unicity problems. So, we shall state the main results of this paper from the following four parts in turn.
A. Uniformization problems
Let be a non-compact Hermitian manifold of complex dimension At first, we generalize the notion of algebroid functions from to by means of some technical treatment, see Definition 2.15. Let be a -valued algebroid function on Assume that is a -leaf meromorphic function element of namely, is a local -valued meromorphic solution of equation (1) in a neighborhood of where is a branching point with order of If is a non-singular branching point of then we describe the asymptotic behavior of as approaches i.e., we show that can be expanded into the Pusieux series (see Theorem 2.14)
with around in a local holomorphic coordinate where are meromorphic functions in Pusieux series plays an important role in construction of the -leaf complex manifold of and so is further uniformized by lifting to That is:
Theorem A1.
Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function on via the natural projection such that
Theorem A1 can be applied directly to the extension of the fields of meromorphic functions. Let be a generic -sheeted analytic covering of We denote by the fields of meromorphic functions on respectively. As a consequence of Theorem A1, we deduce that
Corollary A2.
is a -degree algebraic extension of
B. Nevanlinna theory
Nevanlinna theory in several complex variables has been developed rapidly, due to the outstanding work of a number of excellent scholars. For example, referring to Bott-Chern [8], Carlson-Griffiths-King [10, 20], H. Fujimoto [17], K. Kodaira [25], Noguchi-Winkelmann [31, 32], M. Ru [38], B. Shiffman [40], F. Sakai [42], W. Stoll [48, 49, 52] and Y. T. Siu [53], etc., and also refer the reader to A. Atsuji [1, 2, 3] and X. J. Dong [14, 15], etc.. In this paper, our central work is to establish the Nevanlinna theory of algebroid functions on a non-compact complete Kähler manifold based on estimates of Li-Yau [27] on heat kernels and so we can use which to derive the Picard’s theorem.
In this part and the following Parts B, C, D, is a non-compact complete Kähler manifold. Let be a -valued algebroid function on See definition in Section 5.3, the Nevanlinna’s functions of are defined under two kinds of different curvature conditions. To unify the form, we write
(b1) has non-negative Ricci curvature and maximal volume growth
Refer to [44, 56, 57], we see that such manifolds exist under this curvature condition. Let be the Ricci form of then we obtain:
Theorem B1 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Corollary B2 (Defect relation).
Assume the same conditions as in Theorem B1. Then
We obtain the following Picard’s theorem:
Corollary B3 (Picard Theorem).
Let be a transcendental -valued algebroid function on Then, can omit distinct values at most in
(b2) has non-positive sectional curvature
Let be the curvature function defined by (24) in Section 4.3. Under this curvature condition, we obtain:
Theorem B4 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Corollary B5 (Defect relation).
Assume the same conditions as in Theorem If satisfies the growth condition
then
We obtain the following Picard’s theorem:
Corollary B6 (Picard Theorem).
Let be a nonconstant -valued algebroid function on satisfying the growth condition
Then, can omit distinct values at most in
C. Propagation theory of algebraic dependence
Propagation theory of algebraic dependence was introduced by L. Smiley [47]. Since it has important applications in unicity problems, then it aroused widespread concern. For example, the reader may refer to Y. Aihara [5, 6, 7], Dulock-Ru [12], S. Drouilhet [13], H. Fujimoto [18, 19], S. Ji [23, 24] and W. Stoll [51], etc.. The paper extends this theory to the algebroid functions via the Nevanlinna theory established in Part B.
Let be an effective divisor on where are irreducible hypersurfaces of Let be a positive integer (allowed to be ). Set
Let be hypersurfaces of satisfying that if or if for Let be distinct values in and be positive integers (allowed to be ). Denote by the set of all nonconstant algebroid functions on Set
Let
be the set of all such that
In addition, if has non-positive sectional curvature, then we ask for
Let be the hyperplane line over We denote by the set of all indecomposable hypersurfaces in such that for some where and is the natural projection on the -th factor. Put
with and
Theorem C1.
Let be a -valued algebrod function in for Assume that are -related on for some If then are -related on
Let be a positive integer. Set
Fix any Assume that for at least one Let be the set of all such that on Put
Corollary C2.
If then contains exactly one element.
D. Unicity problems
Five-value theorem is a well-known theorem in unicity theory, which states that two nonconstant meromorphic functions on are identical if they share five distinct values (ignoring multiplicities) in This theorem was later extended to algebroid functions on by G. Valiron [59] and further improved by Sun-Gao [41]. In our study, a sharp five-value type theorem for algebroid functions on a non-compact complete Kähler manifold is proved by applying the propagation theory of algebraic dependence established in Part C.
(d1) has non-negative Ricci curvature and maximal volume growth
Theorem D1.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies
with then
Corollary D2 (Five-value type theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively. If share distinct values ignoring multiplicities in then
(d2) has non-positive sectional curvature
Theorem D3.
Let be nonconstant -valued, -valued algebroid functions on respectively satisfying the growth condition
Let be distinct values in Assume that for all If satisfies
with then
Corollary D4 (Five-value type theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition
If share distinct values ignoring multiplicities in then
2. Algebroid functions on Hermitian manifolds
2.1. Notion of algebroid functions
First of all, we shall illustrate how the notion of algebroid functions on can be generalized to non-compact Hermitian manifolds. This is the starting point of a theory on algebroid functions we would like to build in this paper.
Let be a -dimensional non-compact Hermitian manifold. We consider an irreducible algebraic equation of the unknown function over :
where are holomorphic functions locally defined on such that they define a meromorphic mapping
Let be the indeterminacy locus of which is an analytic set in of complex codimension not less than if or an empty set if So, is a holomorphic mapping of into
Remark 2.1.
The local definition for means that a collection of holomorphic functions i.e., is a holomorphic function for every satisfying that
where is an open covering of Whence, we see that is well defined on and it hence defines a meromorphic function on for
2.1.1. Meromorphic function elements
Set
Then, it gives that
Definition 2.2.
Let A pair is called a regular function element, if is holomorphic in a neighborhood of A pair is called a meromorphic function element, if is meromorphic in a neighborhood of A pair is said to be subordinate to equation if is a local solution of equation in a neighborhood of
Theorem 2.3.
Let be a pair such that
Then there exists a unique regular function element subordinate to equation satisfying
Proof.
The theorem can be showed by employing implicit function theorem. Here, we present an alternative proof of complex analysis. Taking a such that (see Remark 2.1). We write in the form
where
The conditions mean that
By the continuation, there exist such that
for all satisfying and In addition, those are in In further, restrict to then we obtain
Employing Rouché’s theorem, for any fixed with and have the same number of zeros in the disc In fact, the number of zeros is 1, since in has only one zero in the disc when This implies that uniquely determines a function in a smaller geodesic ball neighborhood of with radius such that
Evidently, we see that is continuous on
Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking a local holomorphic coordinate around such that Apply Hartog’s theorem, it is sufficient to show that is holomorphic in every holomorphic direction at For and , set
Note that
as That is,
Thus, is holomorphic in the holomorphic direction at This completes the proof. ∎
Definition 2.4.
A point is called a critical point, if either or has a multiple root as a polynomial in We use to denote the set of all critical points, called the critical set. A regular point means a point in where is called the regular set. Moreover, denote by the set of all points in such that has a multiple root as a polynomial in where is called the multiple set and a point in is called a multiple point.
For any the fundamental theorem of algebra asserts that
has exactly distinct complex roots. Therefore, Theorem 2.3 implies that
Corollary 2.5.
For any there exist exactly distinct regular function elements subordinate to equation
To see the common roots of and we define the resultant of and by a determinant:
where for According to Sylvester’s elimination method, we can see that if and only if or has a multiple root as a polynomial in i.e., This gives that is an analytic set of complex dimension Define the discriminant of by
Assume that splits into linear factors, say
A direct computation leads to
If then is a multiple root of as a polynomial in If we set and rewrite (1) as where
Since then one can see that is a zero of i.e., a pole of In fact, we will deduce from Theorem 2.23 (in Section 2.3) that if and only if To conclude, we see that Corollary 2.5 can be extended to meromorphic function elements as follows:
Corollary 2.6.
For any there exist exactly distinct meromorphic function elements subordinate to equation
2.1.2. Analytic continuation of meromorphic function elements
Recall that a meromorphic function element on is a pair where is a meromorphic function defined in a neighborhood of in As a matter of convenience, we write
for short without confusion. Instead of we sometimes take a geodesic ball neighborhood centered at with radius Let be any two meromorphic function elements on Define an equivalent relation as follows: if and only if
Set
We equip with a topology. A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of Apply analytic continuation, we can regard that
A neighborhood of is defined by a set (containing ) of all such that is a direct analytic continuation of satisfying where is a neighborhood of in Namely, In particular, a -neighborhood of (satisfying that ), is defined by the set (which contains ) of all such that is a direct analytic continuation of satisfying that In other words, it is defined to be A subset is called an open set, if either is an empty set or every point is an inner point, i.e., there exists a -neighborhood of such that It is not hard to check that it defines a topology of
Theorem 2.7.
is a Hausdorff space.
Proof.
Pick If then we can consider -neighborhoods of and of with Evidently, we have If for then we can take a function element It implies that in a small neighborhood of By the uniqueness theorem of analytic functions, we obtain but it contradicts with ∎
Next, we introduce the notion of analytic continuation (by which we mean both holomorphic continuation and meromorphic continuation of meromorphic function elements.
Definition 2.8.
Let be two meromorphic function elements on Let be a curve with and We say that is an analytic continuation of along if
for any there is a meromorphic function element on
for any and any there exists such that for all and that is a direct analytic continuation of
Remark 2.9.
The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Indeed, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve. Thus, the analytic continuation is symmetric.
Theorem 2.10 (Analytic continuation).
Let be a curve. Then, a function element can continue analytically to along and attain to a unique function element
Proof.
Set
Then, we have We show Otherwise, we assume that By Corollary 2.6, there are exactly function elements at say that
Take small enough such that where Then there is a unique integer such that Notice that the domain of definition of is so for any one can redefine function elements Combined with the definition of it yields immediately that can continue analytically to along i.e., is an analytic continuation of along However, it contradicts with the definition of Thus, we obtain The uniqueness of is immediate due to the uniqueness theorem of analytic (meromorphic) functions. ∎
Corollary 2.11.
Let be distinct function elements at respectively. Then can continue analytically to each other along any curve in
Theorem 2.12 (Connectivity).
Any two function elements in can continue analytically to each other along a curve in
Proof.
Let be distinct function elements at every fixed With the aid of Corollary 2.11, then we only need to prove that all can continue analytically to each other. Let be the largest integer such that there are function elements in that can continue analytically to each other along a curve. Clearly, we have Without loss of generality, one may assume that these function elements are So, we can write
where are holomorphic functions in a neighborhood of By Vieta’s theorem, we obtain
It is equivalent to show that: can continue analytically to the whole Apply Theorem 2.10, can continue analytically to along a curve at any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., continue analytically to Hence, holds. If then we are able to determine another set of using the rest function elements However, this implies that equation (1) is reducible, a contradiction. ∎
Corollary 2.13.
is a path-connected space.
2.1.3. Properties of meromorphic function elements near multiple points
We discuss the properties of function elements near multiple points. Let be a pair satisfying that and
Take a sufficiently small neighborhood of such that on By the continuation of there exists a number such that on Let then there exist distinct regular function elements and in respectively. Make use of Theorem 2.10, then for a curve with and are able to continue analytically to along In other words, there exists a permutation of such that can continue analytically to along respectively. By symmetrical property, can also continue analytically to along However, will not, in general, continue analytically back to along respectively. In the following, we will investigate the analytic continuation of a function element along a closed Jordan curve. Let and pick as a closed Jordan curve passing by around Without loss of generality, we extend analytically along started at When returns to we attain to a function If then one continues to extend analytically along Repeat this action, then we will see that cycles on and on periodically. Let be the smallest period such that returns to itself, we obtain We say that is a branching point with order (or multiplicity) of if Repeatedly, then is divided into finite groups. Let the number of groups. If the -th group has members, then we have
Without loss of generality, we consider the first group: Take a proper there is a biholomorphic mapping such that
where stands for the polydisc centered at with polyradius in and is a local holomorphic coordinate of in So, we have
Set and write Next, we shall show that is a single-valued analytic function on where is the unit disc centered at in the -plane. Evidently, is holomorphic on We still need to prove that is a removed singularity of In doing so, it is sufficient to show that is bounded on due to Riemann’s analytic continuation. There exists a number such that on for Combining (1) with on we obtain
on If we obtain Thus, on and this bound does not depend on Whence, is bounded on This means that is holomorphic on and of the Taylor expansion
near Let be the set of all branching points of in Note that the value of is independent of if is sufficiently small. If is a non-singular branching point of i.e., does not lie in a crossing of components of then is smooth whenever is small enough. By and Riemann’s analytic continuation (because the boundedness of on and is properly chosen if necessary), there is a biholomorphic mapping with
such that can be expanded into the Pusieux series
near where are holomorphic functions in satisfying that for
According to the above arguments, we show that is a -valued analytic function near In other words, if we cut along a (real) hypersurface which passes through then is separated into single-valued analytic components, and these components replace each other when they continue analytically along a curve around in Observe the local expansion of is a -multiple -valued branching point (with order ) of
Assume that is not a singular branching point of Set we write
Similarly, we have
If then
Choose such that Put then
Thus, i.e., the case turns to the case Note that is a pole of order of
To conclude, for each we have confirmed that there are distinct single-valued meromorphic function elements subordinate to equation (1) if is not a branching point of If is a branching point of then we can have several multi-valued function elements subordinate to equation (1), say with where gives a branching point of with order In particular, if is a non-singular branching point of then there is the Pusieux expansion
near in a local holomorphic coordinate Note that is a -multiple root of namely, can separate into single-valued components with value at If then can separate into single-valued components which take at In particular, is a meromorphic function element for Furthermore, is a regular function element if also. By analytic continuation, equation (1) uniquely determines a -valued meromorphic function on outside a set of branching points of What is more, is separated into single-valued meromorphic components on a cut-manifold, say where is any real hypersurface of passing by all the branching points of such that is connected. Let denote the set of all the branching points of which is an analytic set contained in called the branching set of We obtain:
Theorem 2.14.
Let be a -valued function element subordinate to equation If is not a singular branching point of then can be expanded into the convergent Pusieux series
in a local holomorphic coordinate near such that and for where is the branching set in a local neighborhood of and are meromorphic functions in
Now, we introduce the definition of algebroid functions on a non-compact Hermitian manifold as follows.
Definition 2.15.
A -valued algebroid function on is defined by an irreducible algebraic equation
where are holomorphic functions locally defined on such that they define a meromorphic mapping from into
Remark 2.16.
By Definition 2.15, we see that the condition for is also necessary due to Vieta’s theorem. In other words, a -valued algebroid function on is defined equivalently by an irreducible algebraic equation
where are meromorphic functions on
2.2. Multi-leaf complex manifolds of algebroid functions
In this subsection, we will show how to construct a class of complex manifolds using algebroid functions. Such a class of manifolds have a remarkable feature of multifoliation. What’s important is that these multi-leaf complex manifolds are a way to the uniformization of algebroid functions. First of all, we deal with a simple situation, i.e., we equip with a complex structure such that is a non-compact complex manifold.
As a matter of convenience, a multi-valued function element is also called the multi-leaf function element; a (non-holomorphic) meromorphic function element is also called the polar function element; the polar function elements and multi-leaf function elements are called the singular function elements.
Let be an arbitrary function element. Let us consider a neighborhood of which is defined by such a set of all such that is a direct analytic continuation of satisfying i.e., we have Choose a proper if necessary, we may have a biholomorphic mapping such that in which is the unit ball centered at 0 in Define a natural mapping
via
Evidently, is a homeomorphism. Its inverse is
Set We show that is a complex atlas of Let It suffices to show the transition mapping
is a biholomorphic mapping. This is automatically a homeomorphism. Note that for each we have in a small neighborhood of i.e., A direct computation gives that
for Similarly, we obtain
for Hence, is a biholomorphic mapping. On the other hand, for any there corresponds to distinct function elements Therefore, we conclude that
Theorem 2.17.
is a non-compact -leaf complex manifold of complex dimension
Set
Corollary 2.18.
is a non-compact -leaf complex manifold of complex dimension which is a submanifold of
In what follows, we shall construct a -leaf complex manifold determined by all functions elements subordinate to equation (1). Let be any two function elements subordinate to equation (1). We define an equivalent relation as follows: if and only if
for some two single-valued components of respectively, where is the branching set of for Now, put the multi-leaf function elements in Set
Remark 2.19.
If then it corresponds to meromorphic function elements in In particular, if then reduce to regular function elements. If then it can correspond to multi-leaf function elements in with where can be separated into single-valued components.
Definition 2.20.
A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branching set of If so, one can regard that
Let A neighborhood of is a set which contains the following elements:
all such that is a direct analytic continuation of satisfying
all satisfying
where is a neighborhood of In particular, a -neighborhood of (satisfying ), is defined by the set which contains the following elements:
all such that is a direct analytic continuation of satisfying
all satisfying
A subset is called an open set, if either is an empty set or every point is an inner point, namely, there exists a -neighborhood of such that If is a -leaf function element, then has leaves. In this case, we call a -leaf neighborhood of This gives a topology of and this topology contains the topology of
Theorem 2.21.
is a connected Hausdorff space.
Proof.
Hausdorff property of can be showed using the similar arguments as in the proof of Theorem 2.7. Next, we prove that is connected. If not, one can write where are two disjoint open subsets of Set It is evident that Notice that are two disjoint open subsets of since is an analytic set of complex codimension This means that is not connected, but it is a contradiction. ∎
Next, let us continue to endow with a complex structure such that is a complex manifold. Take any -leaf function element Assume that is a non-singular branching point of Apply Theorem 2.14 ( is properly chosen if necessary), there exists a biholomorphic mapping such that can be expanded into the Pusieux series
satisfying that and for
Let be a -leaf neighborhood of containing the following elements:
all such that is a direct analytic continuation of satisfying
all satisfying
Let be single-valued connected components of Define a mapping
by
where It is easy to check that is a homeomorphism and its inverse is given by
where is a positive integer such that Let be the set of all function elements such that is a singular branching point of some multi-valued component of and let be the set of all corresponding singular branching points. It is evident that is an analytic set of complex dimension not greater than Set
We will show that is a complex atlas of Let It is sufficient to prove the transition mapping
is a biholomorphic mapping. This is automatically a homeomorphism. Via a similar discussion, one only needs to deal with the situation where is a -leaf singular function element and is just a regular function element. For each there exists a single-valued component of such that in some neighborhood of i.e., It is clear that since all the function elements in are meromorphic. A direct computation leads to
where is a positive integer such that Hence, is a holomorphic mapping. On the other hand, its inverse is
where is the -th single-valued component of Evidently, is also a holomorphic mapping. Hence, is a non-compact -leaf complex manifold of complex dimension
Next, we consider the singular part Let be the natural projection by mapping a function element into Then, one sees that is a generic -sheeted analytic covering and
Since is an analytic set of complex dimension not greater than then the complex structure of can be viewed as a natural extension of the complex structure of This implies that we can naturally extend the complex structure of to via Therefore, we obtain:
Theorem 2.22.
is a non-compact -leaf complex manifold of complex dimension called a complex manifold of a -valued algebroid function on
2.3. Zero and pole divisors of algebroid functions
Let be a -valued algebroid function on defined by equation As we have known in Section 2.1, is a pole of if and only if is a zero of However, it provides nothing about the relationship between the multiplicities of a pole of and a zero of The following theorem shows that they agree. Let be single-valued meromorphic components of As a matter of convenience, we write
Theorem 2.23.
We have
Proof.
Let be a zero of Since is an empty set or an analytic set of complex codimension 2, then we have in the sense of divisors, and so we may assume that It yields that
where denotes the number of -valued points of a function at Without loss of generality, suppose that are all single-valued components such that which take at and suppose that are all single-valued components such that which take at Then, we have By Vieta’s theorem, it yields that
Combined this with (4), we see that reaches the maximal order of a pole of Since then we obtain
It follows that ∎
Theorem 2.24.
We have
Proof.
Set Apply the similar arguments as in the proof of Theorem 2.23, we can prove the theorem. ∎
3. Uniformization of algebroid functions on Hermitian manifolds
This section works on the uniformization of algebroid functions on a non-compact Hermitian manifold. As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing multi-leaf complex manifolds. Our idea is motivated by the work of W. Stoll [50] on algebroid reduction.
3.1. Algebroid reduction
Let be a non-compact complex manifold of complex dimension Let be a generic -sheeted analytic covering, namely, it is a proper and surjective holomorphic mapping with branching set i.e., the set of all points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex dimension It is evident that
is a proper, surjective, local biholomorphic mapping of sheet number and hence a -sheeted analytic covering. By Remmert [39] (see Andreotti-Stoll [4] also), the exceptional set
is an analytic set satisfying that and Moreover, and for Let the mapping degree of at is defined in such manner: if is a relatively compact neighborhood of such that then
which is independent of the choice of Note that and if and only if By W. Stoll [50], for any we have
Let be a complex vector space of complex dimension with a basis Let be the projection of where is the natural projection mapping. Let be the -fold symmetric tensor product of i.e.,
where
for any is called the symmetric tensor product of and is the -order symmetric group. It is evident that is a subspace of -fold tensor product space with Again, put
where for any Naturally, is embedded into by the identity mapping Thus, we obtain
Definition 3.1.
Let be a nonempty open subset of Let be a continuous mapping. Then, the algebroid reduction of via is a mapping defined by
Theorem 3.2 (Stoll, [50]).
is a continuous mapping. If is holomorphic, then is also holomorphic and extends to a holomorphic mapping
Definition 3.3.
Let be a meromorphic mapping with indeterminacy set Then, the algebroid reduction of via is a meromorphic mapping with indeterminacy locus defined by
Theorem 3.4 (Stoll, [50]).
Let be a meromorphic mapping with indeterminacy locus Let be the algebroid reduction of via Fix any Assume that there exist a nonempty neighborhood of such that and a reduced representation If is a representation of on then the algebroid reduction of via is a representation of If is reduced, then is also reduced.
According to Theorem 3.4, we obtain the following commutative diagram:
Theorem 3.5 (Stoll, [50]).
Let be a differentiably non-degenerate and irreducible meromorphic mapping. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via
3.2. Uniformization problems
In what follows, we consider the case that For a convenience, we write and in short. Take a basis of Then for any it has a unique representation
for some numbers It gives a natural isomorphism defined by
which yields that
If then (6) defines an algebraic equation
Let be the maximal integer in such that Then, equation (7) reduces to Let be the roots of the equation. Apply Vieta’s theorem, it is immediate that
So, is decomposable, i.e.,
By this with (5)
To conclude, we obtain
Lemma 3.6.
Let be a differentiably non-degenerate and irreducible meromorphic mapping. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function which separates the fibers of such that is the algebroid reduction of via
Theorem 3.7 (Uniformization theorem).
Let be a -valued algebroid function on Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function such that
Proof.
Let be defined by equation (1). It gives a differentiably non-degenerate and irreducible meromorphic mapping
Set with a basis Since (8), then we can write
in the sense of isomorphisms. Further, the Vieta’s theorem leads to
in the sense of isomorphisms. On the other hand, Lemma 3.6 says that there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of (generic) sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via Hence, we have
with
Let be -valued points of Then
By
we have the theorem proved. ∎
Treat the -leaf complex manifold of a -valued algebroid function on we have the natural projection of (generic) sheet number via which is a proper surjective holomorphic mapping.
Corollary 3.8.
Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function via the natural projection such that
Let be a generic -sheeted analytic covering of a non-compact Hermitian manifold with branching set Denote by (resp. ) the field of all meromorphic functions on (resp. ).
We give the algebraic relation between and as follows.
Corollary 3.9.
is a -degree algebraic extension of
Proof.
For a meromorphic function on we see that is at most a -valued algebroid function with branching set on So, there exist meromorphic functions on such that
It implies that is a finite algebraic extension of of degree ∎
4. Nevanlinna theory of meromorphic functions on Kähler manifolds
Let be a non-compact complete Kähler manifold of complex dimension with Laplace-Beltrami operator associated to Kähler metric Recall that the Kähler form of is defined by
in a local holomorphic coordinate Moreover, we define the Ricci form of by
Invoking Wirtinger’s formula, we see that the Riemannian volume element of is
4.1. First Main Theorem
Let be an arbitrary relatively compact domain with smooth boundary in Fix a point and denote by the positive Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
where is the Dirac function on The harmonic measure on with respect to is defined by
where is the Riemannian volume element of and is the inward normal derivative on
Next, we introduce Nevanlinna’s functions. Let be a meormorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of and here
so that
The characteristic function of for is defined by
Let The proximity function of for with respect to is defined by
where denotes the spherical distance on defined by
with
The counting function of for with respect to is defined by
We need the following Dynkin formula:
Lemma 4.1 (Dynkin formula, [3, 16]).
Let be a -class function on outside a polar set of singularities at most. Assume that Then
We are to show the First Main Theorem:
Theorem 4.2 (First Main Theorem).
Assume that Then
Proof.
Apply Dynkin formula (see Lemma 4.1) to we obtain
Note that the first term on the left hand side equals and the right hand side equals
Hence, we have the theorem proved. ∎
4.2. Second Main Theorem for non-negative Ricci curvature
In the subsection, we assume that has non-negative Ricci curvature and maximal volume growth, i.e.,
where is the volume of geodesic ball centered at a fixed reference point with radius Denote by the Riemannian distance function of from and by the area of which is the boundary of A global Green function of for with a pole at means that
where is the Dirac function on Put
By Li-Tam-Wang [26], there is a number independent of such that
as
If then there exists (see, e.g., [60, 61]) the unique minimal positive global Green function of for with a pole at According to the estimate due to Li-Yau [27] and (10), there exists a constant depending only on such that
Colding-Minicozzi II [11] (see Li-Tam-Wang [26] also) proved the following asymptotic behavior of :
Lemma 4.3.
For we have
Define the relatively compact -domain for by
Then, exhausts whenever is an increasing sequence with as Moreover, the boundary of is a submanifold of for almost every due to Sard’s theorem. Set
It is evident that gives the positive Green function of for with a pole at satisfying Dirichlet boundary condition. Also, denote by the harmonic measure on with respect to
If then curvature condition implies that is a parabolic Riemann surface, so each global Green function for must change sign. Again, since has maximal volume growth, we see that is conformally equivalent to i.e., there exists a biholomorphic mapping such that We equip with the pull-back metric induced from the standard Euclidean metric on Since gives a global Green function of for with a pole at then it corresponds to a global Green function of for with a pole at :
In this case, we define the relatively compact -domain for by
It is clear that
gives the Green function of for with a pole at satisfying Dirichlet boundary condition. A direct computation gives that the harmonic measure on with respect to is
where is the Riemannian volume element of
Let be a meromorphic function. Take in Section 4.1, then we have the Nevanlinna’s functions of :
4.2.1. Estimates of harmonic measures
We shall use the asymptotic behavior of (see Lemma 4.3) to derive both lower and upper bounds of the harmonic measure on
Theorem 4.4 (Gradient estimates of Green functions).
We have
If then
If then there exists a constant such that
holds for all sufficiently large
Proof.
If then is just the geodesic disc centered at with radius It means that for Thus, it follows from (12) that
where is the inward normal derivative on Hence, holds.
Next, we show that holds. Note that is continuous outside and the cut locus of has measure zero, one may assume without loss of generality that Since and as by L’Hôspital’s rule, then it yields from Lemma 4.3 that
for where stands for the inward normal derivative on On the other hand, Lemma 4.3 leads to
Hence, we see that approaches increasingly a geodesic ball centered at with radius as This also implies that
Combining the above, we obtain
Therefore, there exists a constant such that
holds for all sufficiently large ∎
Corollary 4.5 (Estimates of harmonic measures).
We have
If then
If then
for all sufficiently large In the above, is the Riemannian volume element of
Proof.
The corollary follows immediately from the following relation
4.2.2. Calculus Lemma
In order to establish the calculus lemma which plays an important role in the proof of the second main theorem later, we need to estimate the bounds of the Green function for in terms of integral forms. Set
Lemma 4.6 (Estimates of Green functions).
We have
If then
If then for any there exists sufficiently large such that for we have
Proof.
Theorem 4.7 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a positive constant and a subset with finite Lebesgue measure such that
holds for all outside
Proof.
If then is the geodesic disc centered at with radius It yields from Theorem 4.6 that
Set
Then
which yields from Corollary 4.5 that
On the other hand, applying Borel lemma to twice, then for any there exists a subset with finite Lebesgue measure such that
holds for all outside Combining (15) with (16), we have
holds for all outside where is a constant.
Next, we show the case where Invoking Lemma 4.6, for any there exists sufficiently large such that for we have
Notice that as due to Lemma 4.3, then there exists sufficiently large such that for we have
Hence, we conclude that
where
Set
A direct computation leads to
Furthermore, we have
According to Corollary 4.5, there exists sufficiently large such that
holds for all Thus, it follows from (18) that
Using Borel lemma twice, for any , we have
holds for all outside a subset with finite Lebesgue measure. Applying (19) and (20) and noting that as we see that there exists sufficiently large such that for all we have
Therefore, by this with (4.2.2)
holds for all outside where
is a constant. This proves the theorem. ∎
4.2.3. Logarithmic Derivative Lemma
To establish the second main theorem, we still need to prove a logarithmic derivative lemma. Let be the gradient operator on associated with the Kähler metric Let be a meromorphic function on The norm of the gradient of is defined by
in a local holomorphic coordinate where is the inverse of Define
where
Indeed, define for
where
It is not difficult to check that
On take a singular metric
which is satisfied with
To establish the Logarithmic Derivative Lemma, we require the following two lemmas.
Lemma 4.8.
We have
Proof.
A direct computation leads to
which yields from Fubini’s theorem that
Lemma 4.9.
Assume that Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Proof.
Make use of the concavity of , we have
Thus, it yields from Theorem 4.7 that
for all outside The lemma is proved. ∎
Define
Theorem 4.10 (Logarithmic Derivative Lemma).
Let be a nonconstant meromorphic function on Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Proof.
Note that
By Lemma 4.9, for any there exists a subset with finite Lebesgue measure such that
holds for all outside ∎
4.2.4. Second Main Theorem
The characteristic function of is defined by
Theorem 4.11 (Second Main Theorem).
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Proof.
Define a non-negative function on via
where
is a singular volume form on By we obtain
in the sense of currents, in which is the zero divisor of and is the ramification divisor of By Dynkin formula (see Lemma 4.1), it yields that
On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on on and with On each we write
where is a positive smooth function on Given a partition of the unity subordinate to Put
Again, set Then on we have
Fix a we can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
It deduces that
Put Then, we obtain
on Therefore,
on Since is bounded on then it yields that
By this with Dynkin formula (see Lemma 4.1) and Theorem 4.10, for any there is a subset with finite Lebesgue measure such that
holds for all outside By (4.2.4) and (23), then the theorem holds. ∎
Let us consider the defect relation. Define the defect and the simple defect of with respect to respectively by
According to Theorem 4.2, we obtain Since has maximal volume growth, then whenever is nonconstant, as
Corollary 4.12 (Defect relation).
Assume the same conditions as in Theorem Then
4.3. Second Main Theorem for non-positive sectional curvature
In the subsection, we assume that has non-positive sectional curvature. Fix a point Review that denotes the geodesic ball centered at with radius By Sard’s lemma, the boundary of is a submanifold of for almost every Set
where is the pointwise lower bound of Ricci curvature defined by
It remains to use to denote the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and use to denote the harmonic measure on with respect to Let be a meromorphic function. Similarly, the Nevanlinna’s functions are defined by
Refer to Dong [14] (see Atsuji [3] also), we have the Second Main Theorem as follows:
Theorem 4.13 (Second Main Theorem).
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Similarly, we can define the defect and the simple defect of with respect to a value in
Corollary 4.14 (Defect relation).
Assume the same conditions as in Theorem If satisfies the growth condition
then
5. Nevanlinna theory of algebroid functions on Kähler manifolds
Let be a non-compact complete Kähler manifold of complex dimension with Laplace-Beltrami operator associated to its Kähler metric We adopt the same notations such as etc. defined in Section 4.
5.1. First Main Theorem
Let be an arbitrary relatively compact domain with smooth boundary in Fix a point then we have Green function for and harmonic measure on (see definition in Section 4.1).
In what follows, we introduce Nevanlinna’s functions of algbroid functions. Let be a -valued algebroid function with branching set on which is defined by an irreducible algebraic equation
Here, are single-valued meromorphic components of namely, are locally meromorphic on outside Recall that it defines a meromorphic mapping with indeterminacy locus Suppose The characteristic function of for is defined by
where
are called the proximity function, counting function of for respectively. For define
where
In order to unify the form, we write
We show the First Main Theorem:
Theorem 5.1 (First Main Theorem).
Assume that Then
Proof.
With the aid of Vieta’s theorem, we obtain
Using Dykin formula (see Lemma 4.1), then it yields that
From the definition, we see that the left hand side equals
On the other hand, Theorem 2.23 and Theorem 2.24 imply that the right hand side equals
Thus, we obtain
Set then
where
In particular, and Similarly, we have
which means that
Since
and
then we obtain
Put together the above, we conclude that
5.2. Estimates of branching divisors
Let be a -valued algebroid function on We use the same notations such as given in Section 5.1. Our aim is to give an estimate of the branching divisor of Denote by the -valued, -valued components of at any respectively. Then, we have namely, is a -order branching point of for and is hence a -order branching point of The branching divisor of associated to is defined by a divisor on satisfying that
The counting function of branching divisor of for is defined by
where
Recall that the discriminant of is given by
which is a -order determinant, see Section 2.1.
Lemma 5.2.
We have
in the sense of currents.
Proof.
Take any Let be the -valued, -valued components of at respectively, where Since the set of singular branching points of is an analytic set of complex dimension not greater than then one just needs to consider that is a non-singular branching point of By Theorem 2.14, we can write for
in a local holomorphic coordinate near such that has a local coordinate If then is not a pole of One can note from an observation over that there exist terms in which contain the factor i.e., the factor appears at least times in It implies that If then is a pole of By Theorem 2.23, we have
This implies the lemma holds. ∎
Set
Theorem 5.3 (Estimates of branching divisors).
We have
5.3. Second Main Theorem
Let be the -leaf complex manifold of -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.22. Let be the natural projection. Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on for that could be not positive-definite at points in the branching set of Using Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we show that Dykin formula (see Lemma 4.1) still holds under the pullback metric Set Pick a sequence of neighborhoods of such that
Equip with a complete Kähler metric Let Clearly, is a complete Kähler metric on and as Let be the Laplace-Beltrami operators of respectively. Then, weakly as Take a bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and by the harmonic measures of on with respect to respectively. It is evident that as and (weakly) uniformally on as due to
for any Borel subset of Similarly, we also have that (weakly) uniformally on as where denote the volume measures of for respectively. To unify the form, one uses to denote the volume elements of for respectively. Using Dynkin formula, it gives that
for a -class function with on outside a polar set of singularities at most. Hence, it yields immediately by letting that
This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.11 and Theorem 4.13 still hold for under the corresponding curvature conditions, since is an analytic set of complex dimension
5.3.1. with non-negative Ricci curvature
In the subsection, we assume that has non-negative Ricci curvature and maximal volume growth. With the pullback metric has non-negative Ricci curvature and maximal volume growth. Let be the Kähler form of Take where is a bounded domain defined in Section 4.2. Let us recall the Green function for and the harmonic measure on defined in Section 4.2. Then, the characteristic function of is defined by
where
Indeed, define for
where
In order to unify the form, we write
The simple counting function of with respect to is defined by
Assume that We denote by the connected components of centered at respectively, with for Again, define
Then, it is trivial to see that is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and also is the harmonic measure on with respect to Set
where
Define for
and
If is not constant, then it follows from Theorem 4.11 that for distinct values there is a subset with finite Lebesgue measure such that
holds for and all outside
Theorem 5.4 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Proof.
Note that is a -sheeted analytic covering, and two connected components () might meet only at the branching set of and a branching point of is possibly a point in On the other hand, if is a branching point with order of we have multi-leaf function elements subordinate to equation (1), say where with and gives a branching point of with order i.e., So, it is not hard to deduce from the construction of and the definition of that
On the other hand, we have and
Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset with finite Lebesgue measure such that
holds for all outside Hence, we have the theorem proved. ∎
We continue to treat the defect relation. Define the defect and the simple defect of with respect to respectively by
According to Theorem 5.1, we obtain The condition of maximal volume growth gives that if is nonconstant, as
Corollary 5.5 (Defect relation).
Assume the same conditions as in Theorem Then
Corollary 5.6 (Picard Theorem).
Let be a nonconstant -valued algebroid function on Then, can omit distinct values at most in
5.3.2. with non-positive sectional curvature
In the subsection, we assume that has non-positive sectional curvature. Then, also has non-positive sectional curvature under the pullback metric Take where is the geodesic ball centered at with radius Recall the Green function for and the harmonic measure on in Section 4.3. Similarly, the characteristic function of is defined by
where
Indeed, for
where
In order to unify the form, we write
Also, define for
Assume that We denote by the connected components of centered at respectively, here for Again, define
We see that is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and is the harmonic measure on with respect to Put
with
where is the Ricci curvature tensor on Similarly, set
where
Also, define for
If is not constant, then it follows from Theorem 4.13 that for distinct values there is a subset with finite Lebesgue measure such that
holds for and all outside
Apply the similar arguments as in the proof of Theorem 5.4, we obtain:
Theorem 5.7 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Set
Corollary 5.8 (Defect relation).
Assume the same conditions as in Theorem If satisfies the growth condition then
Corollary 5.9 (Picard Theorem).
Let be a nonconstant -valued algebroid function on satisfying the growth condition then can omit distinct values at most in
6. Unicity problems of algebroid functions on Kähler manifolds
In this section, is a non-compact complete Kähler manifold of complex dimension
6.1. Propagation of algebraic dependence
Let be the Fubini-Study form on Given a -valued algebroid function on and Define
Using Dynkin formula (see Lemma 4.1), it is immediate that
On the other hand, since
then we have
Fix any integer A proper algebraic subset of is said to be decomposible, if there exist positive integers with for some integer and algebraic subsets with such that If is not decomposable, then we say that is indecomposable. Given algebroid functions Put
Definition 6.1.
Let be an analytic subset of nonconstant algebroid functions on are said to be algebraically dependent on if there exists a proper indecomposable algebraic subset of such that If so, we say that are -related on
Let be an effective divisor on and fix any positive integer (allowed to be ). Write for distinct irreducible hypersurfaces of and positive integers Define the support of with order at most by
It is clear that Set
Let be hypersurfaces of satisfying that if or if for Let be distinct values in and be positive integers (allowed to be ).
We introduce the notation as follows:
If has non-negative Ricci curvature and maximal volume growth, then we denote by
the set of all such that
If has non-positive sectional curvature, then we denote by
the set of all such that
and
Let stand for the hyperplane line bundle over i.e., is the holomorphic line bundle defined by a point in It gives a holomorphic line bundle over via
where is the natural projection on the -th factor with Let for For , define
with
for where (the first Chern class of ) is a fixed positive -form.
Lemma 6.2.
We have
Proof.
A direct computation leads to
Then, it follows that
Invoking Dynkin formula (see Lemma 4.1), it is therefore
This proves the lemma. ∎
Let be the set of all indecomposable hypersurfaces in such that for some Put
Lemma 6.3.
Assume that and for some Then
Proof.
Lemma 6.4.
Let be a -valued algebroid function in Then
Proof.
By we obtain
in the sense of currents, where and are understood as divisors. Thus, we obtain
By this with Theorem 5.4, we obtain
This proves the lemma. ∎
Set
with
We obtain a criteria for the propagation of algebraic dependence:
Theorem 6.5.
Let be a -valued algebrod function in for Assume that are -related on for some If then are -related on
Proof.
Put
Fix any -valued algebroid function We say that is generic with respect to if for at least one Assume that is generic with respect to Let be the set of all such that on Set
The following is a unicity theorem:
Theorem 6.6.
If then contains exactly one element.
Proof.
Let be a -valued algebroid function in It suffices to show that on Set
Let be the diagonal of Since on then Let be the natural projection on the -th factor with Consider the holomorphic line bundle over where is the hyperplane line bundle over If on then there is a holomorphic section of over such that and Taking then we obtain and On the other hand, implies that
where With the help of Theorem 6.5, we obtain which is a contradiction. This proves the theorem. ∎
6.2. Five-value type theorem
We shall apply the propagation theory of algebraic dependence established in Section 6.1 to the unicity problems of algebroid functions on non-compact complete Kähler manifolds under two kinds of different curvature conditions.
A1 has non-negative Ricci curvature and maximal volume growth
Under this curvature condition, we show that
Theorem 6.7.
Let be transcendental -valued, -valued algebroid functions on respectively. Let be distinct values in Then
Assume for all If then
Assume for all If then
In particular, the conclusions hold for nonconstant if
Proof.
Set and with Then
Making use of Theorem 6.6, we have holds. Similarly, if one sets for then we obtain
which also implies holds. ∎
Corollary 6.8 (Five-value type theorem).
Let be nonconstant -valued algebroid functions on If share distinct values ignoring multiplicities in then
B1 has non-positive sectional curvature
Let be defined by (24). Set
Apply the similar arguments as in the proof of Theorem 6.7, we obtain:
Theorem 6.9.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that satisfy the growth condition Then
Assume for all If then
Assume for all If then
Proof.
The theorem can be confirmed by carrying the same arguments as in the proof of Theorem 6.7. ∎
Corollary 6.10 (Five-value type theorem).
Let be nonconstant -valued algebroid functions on Assume that satisfy the growth condition If share distinct values ignoring multiplicities in then
Remark 6.11.
According to the well-known fact from the case for we see that the number of the shared values is sharp in Corollaries 6.8 and 6.10. However, Theorems 6.7 and 6.9 cannot yield the optimal number of the shared values if In doing so, we need to improve Theorems 6.7 and 6.9 so that one can derive the optimal number of the shared values.
A2 has non-negative Ricci curvature and maximal volume growth
To a general unicity problem for the value sharing ignoring multiplicities, we establish the following theorem:
Theorem 6.12.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies
with then
Proof.
Assume on the contrary that The condition implies that
in the sense of currents, where and are understood as divisors. So, we are led to
where It yields from Lemma 6.3 that
Similarly,
Put together (6.2) and (6.2), it yields from Theorem 5.4 that
Therefore,
But, it contradicts with the assumption of The theorem is proved. ∎
Corollary 6.13.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Then
Assume for all If then
Assume for all If then
Proof.
Corollary 6.13 leads to a sharp five-value type theorem:
Corollary 6.14 (Five-value type theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively. If share distinct values ignoring multiplicities in then
B2 has non-positive sectional curvature
Using the similar arguments as in the proof of Theorem 6.12, we obtain:
Theorem 6.15.
Let be nonconstant -valued, -valued algebroid functions on respectively satisfying the growth condition Let be distinct values in Assume that for all If satisfies
with then
Corollary 6.16.
Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition Let be distinct values in Then
Assume for all If then
Assume for all If then
Corollary 6.16 leads to a sharp five-value type theorem:
Corollary 6.17 (Five-value type theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition If share distinct values ignoring multiplicities in then