### v9

# Nevanlinna Theory of Algebroid Functions on Complete Kähler Manifolds

Abstract.
We generalize the classical Nevanlinna theory for algebroid functions on to a complete Kähler manifold with non-negative Ricci or non-positive sectional curvature. As an application, we give some Picard type theorems and five-value type theorems for algebroid functions under certain conditions.
Key words and phrases:
algebroid functions; Nevanlinna theory; second main theorem; Picard type theorem.2010 Mathematics Subject Classification:
32H30; 32H25Contents
- 1 Introduction
- 2 Algebroid Functions on Hermitian Manifolds
- 3 Uniformization of Algebroid Functions
- 4 Nevanlinna Theory of Meromorphic Functions
- 5 Nevanlinna Theory of Algebroid Functions
- 6 Unicity Problems of Algebroid Functions
1. Introduction
1.1. Motivation
The notion of algebroid functions (see [16, 28]) introduced by H. Poincaré is a generalization of algebraic functions and meromorphic functions. Let us review that a -valued algebroid function on a (nonempty) domain (the complex plane) is defined by an irreducible algebraic equation
where are holomorphic functions with no common zeros on An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G. Darboux deemed that it is an important class of functions. To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a (generic) -sheeted analytic covering of A further fact claims that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
The most concerned problem on algebroid functions seems to be the value distribution. In the 1930s after R. Nevanlinna [22] founded two fundamental theorems concerning the value distribution theory of meromorphic functions in 1925, many famous scholars such as G. Rémoundos, H. Selberg, E. Ullrich and G. Valiron, etc. began to study the value distribution theory of algebroid functions. For example, G. Rémoundos [24] proved a Picard’s theorem which claims that a nonconstant -valued algebroid function on omits distinct values at most in Afterwards, G. Valiron [40] showed without details a five-value type theorem in unicity theory, which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to this topic, including investigations to Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc. More details refer to H. Cartan [7], He-Xiao [16], Niino-Ozawa [19], J. Noguchi [20], M. Ozawa [23], M. Ru [25, 26], Sun-Gao [28], J. Suzuki [29], L. Selberg [30, 31], N. Toda [37, 38] and E. Ullrich [39], etc.
Algebroid function theory in one complex variable didn’t mature until the last 30-40 years. However, there is little literature on the algebroid functions in several complex variables, except for the work of Hu-Yang [15]. Given two nonconstant algebroid functions on a non-compact Hermitian (Kähler) manifold . We are inspired by the following three fundamental problems:
How to realize the uniformization of ?
How many Picard’s exceptional values can have at most?
How many values should share at least to guarantee ?
In this paper, we answer these three problems by generalizing the classical Nevanlinna theory for algebroid functions on to a non-compact complete Kähler manifold with non-negative Ricci or non-positive sectional curvature.
1.2. Main Results
A. Uniformization
Let be a non-compact Hermitian manifold of complex dimension At first, we generalize the notion of algebroid functions from to by means of a technical treatment (see Definition 2.15). Let be a -valued algebroid function on Assume that is a -leaf meromorphic function element of namely, is a local -valued meromorphic solution of equation (7) in a neighborhood of where is a branching point with order of If is a “non-singular” branching point of then we describe the asymptotic behavior of as approaches i.e., we prove that can be expanded into the Pusieux series (see Theorem 2.14)
with around in local holomorphic coordinates where are meromorphic functions in With this Pusieux series, we show that
Theorem A.
Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function on via the natural projection such that
B. Nevanlinna theory
In this part and the following Parts C, D, we utilize to stand for a non-compact complete Kähler manifold. Let be a -valued algebroid function on Refer to the definition in Section 5.3, we have the Nevanlinna’s functions of and on where is the Ricci form of and is a precompact domain containing a fixed reference point To unify the form, we write
The simple defect of with respect to is defined by
(b1) has non-negative Ricci curvature
Assume that is non-parabolic. Set
where denotes the Riemannian volume of the geodesic ball centered at with radius
Theorem B1 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
When we have Hence, Theorem B1 covers the classical second main theorem for with (see Corollary 5.5).
Corollary B2 (Defect Relation).
Assume the same conditions as in Theorem B1. Then
if one of the following conditions is satisfied
satisfies the volume growth condition
where is defined by
is of non-polynomial type growth, i.e., satisfies the growth condition
We have as due to
The condition (2) is relaxed. For example, it holds when with and holds when with or and
Corollary B3 (Picard Type Theorem).
Let be a nonconstant -valued algebroid function on Then, can omit distinct values at most in if one of the conditions in Corollary is satisfied.
(b2) has non-positive sectional curvature
Assume that is non-positively curved satisfying
for constants with that either or where are the sectional curvature and Ricci curvature of respectively. Set
Theorem B4 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
Theorem B4 covers the classical second main theorem for both and the unit ball with Poincaré metric in (see Corollaries 5.7 and 5.8).
Corollary B5 (Defect Relation).
Assume the same conditions as in Theorem B4. If
then
Corollary B6 (Picard Type Theorem).
Let be a nonconstant -valued algebroid function on satisfying Then, can omit distinct values at most in
C. Propagation of algebraic dependence
Propagation problem for algebraic dependence of meromorphic mappings was first studied by L. Smiley [32] and further studied by Y. Aihara [3, 4, 5], S. Drouilhet [9], H. Fujimoto [13, 14], Dulock-Ru [10] and W. Stoll [36], etc.
Assume that is either non-parabolic with non-negative Ricci curvature, or has non-positive sectional curvature satisfying (3).
Let be an effective divisor on where are irreducible hypersurfaces of Let be a positive integer (allowed to be ). Set
Let be hypersurfaces of satisfying that if or if for Let be distinct values in and be positive integers (allowed to be ). Denote by the set of all nonconstant algebroid functions on Set
Let
be the set of all such that
and (2) is satisfied if is non-parabolic with non-negative Ricci curvature, or (5) is satisfied if has non-positive sectional curvature satisfying (3).
Let be the hyperplane line over We denote by the set of all indecomposable hypersurfaces in such that for some where and is the natural projection on the -th factor. Put
with and
Theorem C1.
Let be a -valued algebrod function in for Assume that are -related on for some If then are -related on
Let be a positive integer. Set
Fix any Assume that for at least one Let be the set of all such that on Put
Corollary C2.
If then contains exactly one element.
D. Unicity problems
(d1) has non-negative Ricci curvature
Assume that is non-parabolic satisfying (2). Let be a positive integer. Consider the following condition
with
Theorem D1.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies then
Corollary D2 (Five-Value Type Theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively. If share distinct values ignoring multiplicities in then
(d2) has non-positive sectional curvature
Assume that is non-positively curved satisfying (3).
Theorem D3.
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Assume that for all If satisfies then
Corollary D4 (Five-Value Type Theorem).
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. If share distinct values ignoring multiplicities in then
2. Algebroid Functions on Hermitian Manifolds
2.1. Notion of Algebroid Functions
First of all, we would like to extend the notion of algebroid functions to a non-compact Hermitian manifold. Let be a -dimensional non-compact Hermitian manifold. Consider an irreducible algebraic equation in on :
where are holomorphic functions locally defined on such that they well define a meromorphic mapping
Let be the indeterminacy locus of which is an analytic set in of complex codimension not less than if or an empty set if So, is a holomorphic mapping of into
Remark 2.1.
The local definition for means that a collection of holomorphic functions i.e., is a holomorphic function for all satisfying that
where is an open covering of Whence, we see that is well defined on and it hence defines a meromorphic function on for
2.1.1. Meromorphic function elements
Set
It gives that
Definition 2.2.
Let A pair is called a regular function element, if is holomorphic in a neighborhood of A pair is called a meromorphic function element, if is meromorphic in a neighborhood of A pair is said to be subordinate to equation if is a local solution of equation in a neighborhood of
Theorem 2.3.
Let be a pair such that
Then there exists a unique regular function element subordinate to equation satisfying
Proof.
The theorem can be showed by employing implicit function theorem. Here, we present an alternative proof of complex analysis. Taking a such that (see Remark 2.1). We write in the form
where
The conditions mean that
By the continuation, there exist such that
for all satisfying and In addition, those are in In further, restrict to then we obtain
Employing Rouché’s theorem, for any fixed with and have the same number of zeros in the disc In fact, the number of zeros is 1, since in has only one zero in the disc when This implies that uniquely determines a function in a smaller geodesic ball neighborhood of with radius such that
Evidently, is continuous on
Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking local holomorphic coordinates around such that Apply Hartog’s theorem, it is sufficient to show that is holomorphic in every holomorphic direction at For and , set
Note that
as That is,
Thus, is holomorphic in the holomorphic direction at This completes the proof. ∎
Definition 2.4.
A point is called a critical point, if either or has a multiple root as a polynomial in We use to denote the set of all critical points, called the critical set. A regular point means a point in where is called the regular set. Moreover, denote by the set of all points in such that has a multiple root as a polynomial in where is called the multiple set and a point in is called a multiple point.
For any the fundamental theorem of algebra asserts that
has exactly distinct complex roots. Therefore, Theorem 2.3 implies that
Corollary 2.5.
For any there exist exactly distinct regular function elements subordinate to equation
To see the common roots of and we define the resultant of and by a -order determinant:
where for According to Sylvester’s elimination method, we can see that if and only if or has a multiple root as a polynomial in i.e., This gives that is an analytic set of complex dimension Define the discriminant of by
Assume that splits into linear factors, say
A direct computation leads to
If then has a multiple root as a polynomial in If we set and rewrite (7) as where
Since then one can see that is a zero of i.e., a pole of In fact, we will deduce from Theorem 2.22 (in Section 2.3) that if and only if To conclude, we see that Corollary 2.5 can be extended to meromorphic function elements as follows:
Corollary 2.6.
For any there exist exactly distinct meromorphic function elements subordinate to equation
2.1.2. Analytic continuation of meromorphic function elements
Recall that a meromorphic function element on is a pair where is a meromorphic function defined in a neighborhood of in As a matter of convenience, we write
for short without confusion. Instead of we sometimes take a geodesic ball neighborhood centered at with radius Let be any two meromorphic function elements on Define an equivalent relation as follows: if and only if
Set
We equip with a topology. A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of By the analytic continuation, we can regard that
A neighborhood of is defined by a set (containing ) of all such that is a direct analytic continuation of satisfying where is a neighborhood of in Namely, In particular, a -neighborhood of (satisfying that ), is defined by the set (which contains ) of all such that is a direct analytic continuation of satisfying that In other words, it is defined to be A subset is called an open set, if either is an empty set or every point is an inner point, i.e., there exists a -neighborhood of such that It is not hard to check that it defines a topology of
Theorem 2.7.
is a Hausdorff space.
Proof.
Pick If then we can consider -neighborhoods of and of where Evidently, we have If for then we can take a function element It implies that in a small neighborhood of By the uniqueness theorem of analytic functions, we obtain but it contradicts with ∎
Next, we introduce the notion of analytic continuation (by which we mean both holomorphic continuation and meromorphic continuation) of meromorphic function elements.
Definition 2.8.
Let be two meromorphic function elements on Let be a curve with and We say that is an analytic continuation of along if
for any there is a meromorphic function element on
for any and any there exists such that for all and that is a direct analytic continuation of
Remark 2.9.
The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Moreover, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve. Thus, the analytic continuation is symmetric.
Theorem 2.10 (Analytic continuation).
Let be a curve. Then, a function element can continue analytically to along and attain to a unique function element
Proof.
Set
Then, we have We show Otherwise, we assume that By Corollary 2.6, there are exactly function elements at say that
Take small enough such that where Then there is a unique integer such that Notice that the domain of definition of is so for any one can redefine function elements Combined with the definition of it yields immediately that can continue analytically to along i.e., is an analytic continuation of along However, it contradicts with the definition of Thus, we obtain The uniqueness of is immediate due to the uniqueness theorem of analytic (meromorphic) functions. ∎
Corollary 2.11.
Let be distinct function elements at respectively. Then can continue analytically to each other along any curve in
Theorem 2.12 (Connectivity).
Any two function elements in can continue analytically to each other along a curve in
Proof.
Let be distinct function elements at every fixed With the aid of Corollary 2.11, then we only need to prove that all can continue analytically to each other. Let be the largest integer such that there are function elements in which can continue analytically to each other along a curve. Clearly, we have Without loss of generality, we shall assume that these function elements are So, we can write
where are holomorphic functions in a neighborhood of By Vieta’s theorem, we obtain
It is equivalent to show that: can continue analytically to the whole Apply Theorem 2.10, can continue analytically to along a curve at any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., continue analytically to Hence, holds. If then we are able to determine another set via the rest function elements However, it implies that equation (7) is reducible, which is a contradiction. ∎
Corollary 2.13.
is a path-connected space.
2.1.3. Properties of meromorphic function elements near multiple points
We discuss the local properties of function elements near multiple points. Let be a pair satisfying that and
Taking a sufficiently small neighborhood of such that on By the continuation of there exists a number such that on Let then there exist distinct regular function elements and in respectively. By the use of Theorem 2.10, then for a curve with and are able to continue analytically to along In other words, there exists a permutation of such that can continue analytically to along respectively. By symmetrical property, can also continue analytically to along However, will not, in general, continue analytically back to along respectively. In the following, we discuss the analytic continuation of a function element along a closed Jordan curve surrounding a multiple point. Take Pick as a closed Jordan curve passing by around Without loss of generality, we only extend analytically along starting from When returns to from along one will attain to a function If then one continues to extend analytically along Repeating this action, we see finally that cycles on and on periodically. Let be the smallest period such that returns to itself, we obtain We say that is a branching point with order (multiplicity) of if Repeatedly, then it is not hard to verify that can be divided into several groups. Let be the number of groups. If the -th group has members, then we have
Without loss of generality, we consider the first group: Take a proper we have a biholomorphic mapping such that
where stands for the polydisc centered at with polyradius in and is a local holomorphic coordinate of in So, we have
Set and write Next, we shall show that is a single-valued analytic function on where is the unit disc centered at in the -plane. Evidently, is holomorphic on We still need to prove that is a removed singularity of In doing so, it is sufficient to show that is bounded on due to Riemann’s analytic continuation. There exists a number such that on for Combining (7) with that on we obtain
on If we obtain Thus, on and this bound does not depend on Whence, is bounded on This means that is holomorphic on and of the Taylor expansion
near Let be the set of all branching points of in Note that the value of is independent of if is sufficiently small. If is a non-singular branching point of i.e., does not lie in a crossing of components of then is smooth whenever is small enough. By and Riemann’s analytic continuation (because the boundedness of on and is properly chosen if necessary), there is a biholomorphic mapping with
such that can be expanded into the Pusieux series
near where are holomorphic functions in satisfying that for
According to the above arguments, we show that is a -valued analytic function near In other words, if we cut along a (real) hypersurface which passes through then is separated into single-valued analytic components, and these components replace each other when they continue analytically along a closed curve around in By a local expansion of is a -multiple -valued branching point (with order ) of
Assume that is not a singular branching point of Set and write
Similarly, we have
If then
Choose such that Put then
Thus, i.e., the case turns to the case Note that is a pole with order of
To conclude, for each we have confirmed that there are distinct (single-valued) meromorphic function elements subordinate to equation (7), if is not a branching point. If is a branching point, then one shows that there are several multi-valued function elements (which contain single-valued function elements) subordinate to equation (7), say where is a -order branching point of satisfying At this time, we call a -order branching point of If is a non-singular branching point of then we have the convergent Pusieux expansion
near in a local holomorphic coordinate Note that is a -multiple root of namely, can separate into single-valued components with value at If then can separate into single-valued components which take at In particular, is a meromorphic function element for Furthermore, is a regular function element if also. By the analytic continuation, equation (7) uniquely defines a -valued meromorphic function on outside a set of branching points of and is separated into single-valued meromorphic components on Let denote the set of all branching points of which is an analytic set contained in called the branching set of As a result, we conclude that
Theorem 2.14.
Let be a -valued function element subordinate to equation If is not a singular branching point of then can be expanded into the convergent Pusieux series
in a local holomorphic coordinate near such that =0 and for where is the branching set in a local neighborhood of and are meromorphic functions in
We conclude the definition of algebroid functions on as follows.
Definition 2.15.
A -valued algebroid function on is defined by an irreducible algebraic equation
where are holomorphic functions locally defined on such that they well define a meromorphic mapping from into
Let denote distinct (single-valued) meromorphic components of Given any -leaf function element of with If denote the single-valued components of then there exists a permutation of such that on In other words, defines a -leaf function element on Thus, any function element of can be represented locally by one or several single-valued components of
2.2. Multi-leaf Complex Manifolds of Algebroid Functions
Let be a function element. A neighborhood of is defined as a set of all such that is a direct analytic continuation of satisfying i.e., Taking a proper if necessary, there exists a biholomorphic mapping such that where is the unit ball centered at the coordinate origin in Define a natural mapping
via
Evidently, is a homeomorphism. Its inverse is
Set We show that is a complex atlas of Let It suffices to show the transition mapping
is a biholomorphic mapping. This is automatically a homeomorphism. Note that for each we have in a small neighborhood of i.e., A direct computation gives that
for Similarly, we obtain
for Hence, is a biholomorphic mapping. On the other hand, for any there corresponds to distinct function elements Therefore, we conclude that
Theorem 2.16.
is a non-compact -leaf complex manifold of complex dimension
Set
Corollary 2.17.
is a non-compact -leaf complex manifold of complex dimension which is a submanifold of
In what follows, we shall construct a -leaf complex manifold determined by all functions elements subordinate to equation (7). Let be any two function elements subordinate to equation (7). We define an equivalent relation as follows: if and only if
for some two single-valued components of respectively, where is the branching set of for Set
Remark 2.18.
If then it corresponds to meromorphic function elements in In particular, if then reduce to regular function elements. If then it can correspond to multi-leaf function elements in where is a -order branching point of satisfying In other words, can be separated into single-valued components for
Definition 2.19.
A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branching set of If so, one can regard that
Let A neighborhood of is a set which contains the following elements:
all such that is a direct analytic continuation of satisfying
all satisfying
where is a neighborhood of In particular, a -neighborhood of (satisfying ), is defined by the set which contains the following elements:
all such that is a direct analytic continuation of satisfying
all satisfying
A subset is called an open set, if either is an empty set or every point is an inner point, namely, there exists a -neighborhood of such that If is a -leaf function element, then has leaves. In this case, we call a -leaf neighborhood of This gives a topology of and this topology contains the topology of
Theorem 2.20.
is a connected Hausdorff space.
Proof.
Hausdorff property of can be showed using the similar arguments as in the proof of Theorem 4.3. Next, we prove that is connected. If not, one can write where are two disjoint open subsets of Set It is evident that Notice that are two disjoint open subsets of since is an analytic set of complex codimension This means that is not connected, but it is a contradiction. ∎
Next, let us continue to endow with a complex structure such that is a complex manifold. Take any -leaf function element Assume that is a non-singular branching point of Apply Theorem 2.14 ( is properly chosen if necessary), there exists a biholomorphic mapping such that can be expanded into the Pusieux series
satisfying that and for
Let be a -leaf neighborhood of containing the following elements:
all such that is a direct analytic continuation of satisfying
all satisfying
Let be single-valued connected components of Define a mapping by
where Note that is a homeomorphism and its inverse is
where is a positive integer such that Let be the set of all function elements such that is a singular branching point of some multi-valued component of and let be the set of all corresponding singular branching points. It is evident that is an analytic set of complex dimension not greater than Set
We will show that is a complex atlas of Suppose It is sufficient to prove the transition mapping
is a biholomorphic mapping. This is automatically a homeomorphism. Via a similar discussion, one only needs to deal with the situation where is a -leaf meromorphic function element and is a regular function element. For every there exists a single-valued component of such that in some neighborhood of i.e., It is clear that because all function elements in are meromorphic. A direct computation leads to
where is a positive integer such that Hence, is a holomorphic mapping. On the other hand, its inverse is
where is the -th single-valued component of Evidently, is also a holomorphic mapping. Hence, is a non-compact -leaf complex manifold of complex dimension
Finally, we treat the singular part Let be the natural projection by mapping a function element into Then, we note that is a generic -sheeted analytic covering such that
Since is an analytic set of complex dimension not greater than then the complex structure of can be viewed as a natural extension of the complex structure of This implies that we can naturally extend the complex structure of to via To conclude, we obtain:
Theorem 2.21.
is a non-compact -leaf complex manifold of complex dimension which is called the complex manifold of
2.3. Zero and Pole Divisors of Algebroid Functions
Let be a -valued algebroid function on defined by equation Let be single-valued meromorphic components of As a matter of convenience, we write
Theorem 2.22.
We have
Proof.
Let be a zero of Since is an empty set or an analytic set of complex codimension 2, then we have in the sense of divisors. Thus, we assume that It yields that
where denotes the number of -valued points of a function at Without loss of generality, suppose that are all single-valued components such that which take at and suppose that are all single-valued components such that which take at Then, we have By Vieta’s theorem, it yields that
Combining this with (10), we see that reaches the maximal order of a pole of Since then we obtain
which follows that ∎
Theorem 2.23.
We have
Proof.
Set Apply the similar arguments as in the proof of Theorem 2.22, we can prove the theorem. ∎
3. Uniformization of Algebroid Functions
This section works on the uniformization of algebroid functions on a non-compact Hermitian manifold. As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing a multi-leaf complex manifold. Our main idea comes from the excellent work of W. Stoll [35] on algebroid reduction.
3.1. Algebroid Reduction
Let be a non-compact complex manifold of complex dimension Let be a generic -sheeted analytic covering, namely, it is a proper and surjective holomorphic mapping with branching set i.e., the set of all points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex dimension It is evident that
is a proper, surjective, local biholomorphic mapping of sheet number and hence a -sheeted analytic covering. By Remmert [27] (see Andreotti-Stoll [2] also), the exceptional set
is an analytic set satisfying that and Moreover, and for Let the mapping degree of at is defined in such manner: if is a relatively compact neighborhood of such that then
which is independent of the choice of Note that and if and only if By W. Stoll [35], for any we have
Let be a complex vector space of complex dimension with a basis Let be the projection of where is the natural projection mapping. Let be the -fold symmetric tensor product of i.e.,
where
for any is called the symmetric tensor product of and is the -order symmetric group. It is evident that is a subspace of -fold tensor product space with Again, put
where for any Naturally, is embedded into by the identity mapping Thus, we obtain
Definition 3.1.
Let be a nonempty open subset of Let be a continuous mapping. Then, the algebroid reduction of via is a mapping defined by
Theorem 3.2 (Stoll, [35]).
is a continuous mapping. If is holomorphic, then is also holomorphic and extends to a holomorphic mapping
Definition 3.3.
Let be a meromorphic mapping with indeterminacy set Then, the algebroid reduction of via is a meromorphic mapping with indeterminacy locus defined by
Theorem 3.4 (Stoll, [35]).
Let be a meromorphic mapping with indeterminacy locus Let be the algebroid reduction of via Fix any Assume that there exist a nonempty neighborhood of such that and a reduced representation If is a representation of on then the algebroid reduction of via is a representation of If is reduced, then is also reduced.
According to Theorem 3.4, we obtain the following commutative diagram:
Theorem 3.5 (Stoll, [35]).
Let be a differentiably non-degenerate and irreducible meromorphic mapping. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via
3.2. Uniformization
In what follows, we consider the case that For a convenience, we write and in short. Take a basis of Then for any it has a unique representation
for some numbers It gives a natural isomorphism defined by
which yields that If then (12) defines an algebraic equation
Let be the maximal integer in such that Then, equation (13) reduces to Let be the roots of the equation. Apply Vieta’s theorem, it is immediate that
So, is decomposable, i.e.,
By this with (11), we have To conclude, we obtain
Lemma 3.6.
Let be a differentiably non-degenerate and irreducible meromorphic mapping. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function which separates the fibers of such that is the algebroid reduction of via
Theorem 3.7 (Uniformization theorem).
Let be a -valued algebroid function on Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function such that
Proof.
Let be defined by equation (7). It gives a differentiably non-degenerate and irreducible meromorphic mapping
Set with a basis By (14), we can write
in the sense of isomorphisms. Further, the Vieta’s theorem leads to
in the sense of isomorphisms. On the other hand, Lemma 3.6 says that there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via Hence, we have
with
Let be -valued points of Then
By
we have the theorem proved. ∎
Treat the -leaf complex manifold of a -valued algebroid function on we have the natural projection of generic sheet number via which is a proper surjective holomorphic mapping.
Corollary 3.8.
Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function via the natural projection such that
4. Nevanlinna Theory of Meromorphic Functions
Let be a non-compact complete Kähler manifold of with Laplace-Beltrami operator The Kähler form of is defined by
in local holomorphic coordinates Moreover, we define the Ricci form of by
4.1. First Main Theorem
Let be an arbitrary relatively compact domain with smooth boundary in Fix a point and denote by the positive Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
where is the Dirac’s delta function on with a pole at The harmonic measure on with respect to is defined by
where is the Riemannian volume element of and is the inward normal derivative on
Next, we introduce Nevanlinna’s functions. Let be a meormorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and
The characteristic function of for is defined by
Let The proximity function of for with respect to is defined by
where denotes the spherical distance on defined by
with
The counting function of for with respect to is defined by
Lemma 4.1 (Dynkin formula, [1, 11, 12]).
Let be a -class function on outside a polar set of singularities at most. Assume that Then
Theorem 4.2 (First Main Theorem).
Assume that Then
Proof.
Apply Dynkin formula (see Lemma 4.1) to we obtain
Note that the first term on the left hand side equals and the right hand side equals
4.2. Second Main Theorem for Non-negative Ricci Curvature
Assume that is non-parabolic with non-negative Ricci Curvature. Fix a reference point and denote by the geodesic ball centered at with radius in Denote by the Riemannian volume of The non-parabolicity of means that
Thus, we have the unique minimal positive global Green function of for which can be written as
where is the heat kernel of on Let be the Riemannian distance function of from By Li-Yau’s estimate [18], there exist constants such that
holds for all
4.2.1. Construction of and Nevanlinna’s functions
For define
Since and then we conclude that is a precompact domain containing satisfying that and In fact, the family exhausts i.e., for any sequence with we have
So, the boundary of can be formulated as
By Sard’s theorem, is a submanifold of for almost all
Set
Note that defines the Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
Let denote the harmonic measure on with respect to defined by
where is the inward normal derivative on is the Riemannian area element of
Let be a meromorphic function. Take in Section 4.1, we have the Nevanlinna’s functions of :
4.2.2. Gradient estimates of Green functions
Lemma 4.3.
For we have
holds for all where is given by
Proof.
According to the definition of Green function for it is immediate that for
Since for then we obtain
Let denote the gradient operator on any Riemannian manifold. Cheng-Yau [8] proved the following theorem.
Lemma 4.4.
Let be a complete Riemannian manifold. Let be a geodesic ball centered at with radius Then, there exists a constant depending only on the dimension of such that
holds for any non-negative harmonic function on where is the Riemannian distance between and and is the lower bound of Ricci curvature of
Theorem 4.5.
There exists a constant such that
Proof.
By
we obtain an estimate for upper bounds of as follows.
Corollary 4.6.
There exists a constant such that
where is the Riemannian area element of
4.2.3. Calculus Lemma and Logarithmic Derivative Lemma
Lemma 4.7 (Borel Lemma).
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
holds for all outside
Proof.
The conclusion is clearly true for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit exists or equals If then Set
Note that exists for almost all Then, we have
This completes the proof. ∎
Set
Theorem 4.8 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then there exists a constant such that for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Invoking Lemma 4.3, we have
Set
A simple computation leads to
In further, we have
Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to then we deduce that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.6 implies that there exists a constant such that
Set Combining the above, we have
holds for all outside ∎
In the following, we establish a logarithmic derivative lemma. Let be a meromorphic function on The norm of the gradient of is defined by
in local holomorphic coordinates here is the inverse of Define the Nevanlinna’s characteristic function of by
where
It is not difficult to show that
On one puts a singular metric
which leads to
Lemma 4.9.
We have
Proof.
Note that
Whence, it concludes from Fubini’s theorem that
Lemma 4.10.
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Define
Theorem 4.11 (Logarithmic Derivative Lemma).
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Note that
By Lemma 4.18, we have for any there exists a subset of finite Lebesgue measure such that
holds for all outside ∎
4.2.4. Second Main Theorem
The characteristic function of is defined by
Theorem 4.12 (Second Main Theorem).
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Define a non-negative function on via
where
is a singular volume form on By we obtain
in the sense of currents, in which is the zero divisor of and is the ramification divisor of By Dynkin formula (see Lemma 4.1)
On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on on and
for On each write
where is a positive smooth function on Given a partition of the unity subordinate to Put
Again, set Then on we have
For any we can take local holomorphic coordinates near and a local holomorphic coordinate near such that
It deduces that
Put Then, we obtain
on Therefore,
on Since is bounded on then it yields that
By this with Dynkin formula and Theorem 4.11, for any there exists a subset of finite Lebesgue measure such that
holds for all outside where one used the fact that Combining (4.2.4) and (19), we have the theorem proved. ∎
4.3. Second Main Theorem for Non-positive Sectional Curvature
Assume that has non-positive sectional curvature satisfying (3). From the definition of curvatures, it is not difficult to see that In particular, we have if has constant sectional curvature Fix a reference point We may assume without loss of generality that is a pole of for a technical reason if necessary, since, otherwise, one can consider the universal covering As long as is equipped with the pull-back metric induced from via then has the same curvatures as ones of In fact, is a Cartan-Hadamard manifold. In the following, we denote by the Riemannian distance function of from on
4.3.1. Construction of and Nevanlinna’s functions
A. Setups for
Due to the curvature conditions, we have a unique minimal positive global Green function of with a pole at for Consider the following initial value problems on :
and
Evidently, the two ODEs have the unique solutions
Using the estimates of we have (see [17, 34])
where
Here, is the standard Euclidean area of For define
Since as and then we deduce that is a precompact domain containing in which exhausts So, the boundary of can be formulated as
Set
which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. Denote by the harmonic measure on with respect to i.e.,
where is the inward normal derivative on is the Riemannian area element of In particular, if In this case, we have
which is just the geodesic ball centered at with radius in Then
Let be the geodesic ball centered at with radius in We give the relation between and as follows.
Set
Theorem 4.13.
For we have and
Hence, there exists a constant independent of such that
Proof.
B. Setups for
We shall extend the settings for to the case that For the function does not make sense. However, since
we can extend continuously to defined by (4).
Notice that is parabolic when i.e., there exist no positive global Green functions for In other words, for Hence, one cannot use to define the domain in this case. However, we note (21) for it is thus natural to define
for Let it then yields from (22) that
To conclude, has the unified form
for In further, we have
for
Let be a meromorphic function. Similarly, one can defined the Nevanlinna’s functions and of as well as on as in Section 4.2.1 and Section 4.2.4.
4.3.2. Calculus Lemma and Logarithmic Derivative Lemma
Lemma 4.14.
For we have for
Proof.
When is the geodesic ball centered at with radius The conclusion holds immediately due to (23). In the following, we consider the other case. Since on then for
This completes the proof. ∎
Theorem 4.15.
There exists a constant such that
holds for
Proof.
When is a geodesic ball centered at with radius Hence, it yields from (4) and (23) that
This implies the conclusion holds. For by Lemma 4.4 with (3), (4) and (20) (see Remark 5 in [34] also)
for some large constant which depends only on the dimension By (20) again
which gives for So, we conclude that
where ∎
Corollary 4.16.
There exists a constant such that
holds for where is the Riemannian area element of
Set
Theorem 4.17 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then for any there exists a set of finite Lebesgue measure such that
holds for outside where is defined by
Proof.
Lemma 4.14 yields that
Differentiating to obtain
In further
Applying Borel’s lemma to the left hand side on the above equality twice, then we see that there exists a set of finite Lebesgue measure such that
holds for outside Meanwhile, Corollary 4.16 means that there exists a constant such that
Substituting (27) into (4.3.2), we conclude that
for outside , where ∎
Lemma 4.18.
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for outside where is defined by
Proof.
Theorem 4.19 (Logarithmic Derivative Lemma).
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for outside where is defined by
Proof.
4.3.3. Second Main Theorem
Theorem 4.20 (Second Main Theorem).
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
Proof.
Applying the similar arguments as in the proof of Theorem 4.12 and combining Theorem 4.17 with Theorem 4.19, we conclude that for any there exists a subset of finite Lebesgue measure such that
holds for all outside
In the following, we estimate When we have
This implies that the theorem holds. We consider the other case that Since as we deduce that
Thus, we conclude that
This completes the proof. ∎
5. Nevanlinna Theory of Algebroid Functions
Let be a non-compact complete Kähler manifold of with Laplace-Beltrami operator associated to the Kähler metric We utilize the same notations such as etc. defined in Section 4.
5.1. First Main Theorem
Let be an arbitrary relatively compact domain with smooth boundary in Fix a point then we have Green function for and harmonic measure on (see definitions in Section 4.1).
In what follows, we introduce Nevanlinna’s functions of algbroid functions. Let be a -valued algebroid function with branching set on which is defined by an irreducible algebraic equation
Here, are single-valued meromorphic components of namely, are locally meromorphic on outside Recall that it defines a meromorphic mapping with indeterminacy locus We can choose such that
The characteristic function of for is defined by
where
are called the proximity function, counting function of for respectively. For define
where
In order to unify the form, we write
Theorem 5.1 (First Main Theorem).
Assume that Then
Proof.
With the aid of Vieta’s theorem, we obtain Dynkin formula yields that
By definition, the left hand side equals
On the other hand, Theorem 2.22 and Theorem 2.23 imply that the right hand side equals
We obtain
Set then
with
In particular, and Similarly, we have
which means that
Since
then we obtain
Put together the above, we conclude that
5.2. Estimates of Branching Divisors
Let be a -valued algebroid function on We use the same notations such as given in Section 5.1. Our aim is to give an estimate of the branching divisor of Denote by the -valued, -valued components of at any respectively. Then, we have namely, is a -order branching point of for and is hence a -order branching point of The branching divisor of associated to is defined by a divisor on satisfying that
The counting function of branching divisor of for is defined by
where
Review that the discriminant of is given by
Lemma 5.2.
We have
in the sense of currents.
Proof.
Take any Let be the -valued, -valued components of at respectively, where Since the set of singular branching points of is an analytic set of complex dimension not greater than then one just needs to consider that is a non-singular branching point of By Theorem 2.14, we can write for
in a local holomorphic coordinate near such that has a local coordinate If then is not a pole of So, one can derive from an observation over that there exist terms in which contain the factor i.e., the factor appears at least times in It implies that If then is a pole of By Theorem 2.22, we have
This implies the lemma holds. ∎
Set
Theorem 5.3 (Estimates of branching divisors).
We have
5.3. Second Main Theorem
Let be the -leaf complex manifold of a -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.21. Let be the natural projection. Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on because could be not positive-definite at points in the branching set of Applying Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we explain that Dynkin formula (see Lemma 4.1) still holds under the pullback metric
Set Picking a sequence of neighborhoods of such that
Let be a partition of the unity subordinate to the open covering of Then, and Fix a complete Kähler metric on Let Evidently, is a complete Kähler metric on with on Since then we may assume that (otherwise, a subsequence can be chosen so that as ). Let be the Laplace-Beltrami operators of on respectively. We have weakly as Take an arbitrary bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and again by the harmonic measures of on with respect to respectively. It is not difficult to show that as and (weakly) uniformally on as due to
for any Borel subset of Similarly, we also have that (weakly) uniformally on as where denote the volume measures of for respectively. To unify the form, one uses to stand for the volume elements of for respectively. Apply Dynkin formula, it gives
for a -class function with on outside a polar set of singularities at most. Hence, it yields by letting that
This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.12 and Theorem 4.20 also hold for under the corresponding curvature conditions, since is an analytic set of complex dimension
5.3.1. with non-negative Ricci curvature
Assume that is non-parabolic. With the pullback metric is also non-parabolic, with non-negative Ricci curvature. Let be the Kähler form of Take where is a bounded domain defined in Section 4.2.1. Recall the Green function for and the harmonic measure on defined in Section 4.2.1. Then, the characteristic function of is defined by
where
Indeed, define for
where
In order to unify the form, we write
The simple counting function of with respect to is defined by
Assume that We denote by the connected components of centered at respectively, with for Again, define
Then, it is trivial to see that is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and also is the harmonic measure on with respect to Set
where
Define for
and
If is not constant, then it follows from Theorem 4.12 that for distinct values there is a subset of finite Lebesgue measure such that
holds for and all outside
Theorem 5.4 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Note that is a -sheeted analytic covering, and two connected components () might meet only at the branching set of and a branching point of is possibly a point in On the other hand, if is a branching point with order of we have multi-leaf function elements subordinate to equation (7), say where with and is a branching point of with order i.e., So, it is not hard to deduce from the construction of and the definition of that
On the other hand, we have and
Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset of finite Lebesgue measure such that
holds for all outside Hence, we have the theorem proved. ∎
When we have By it is trivial to deduce Hence, we obtain:
Corollary 5.5.
Let be a nonconstant -valued algebroid function on with Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
5.3.2. with non-positive sectional curvature
Assume that has non-positive sectional curvature satisfying (3). Note that the curvatures of also satisfy (3) under the pullback metric In the same way, we obtain notations and by setting Apply the similar arguments as in the proof of Theorem 5.4, we can show that
Theorem 5.6 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
If then and Hence, we conclude that
Corollary 5.7.
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Let be the unit ball centered at the origin 0 in Consider equipped with the Poincaré metric
in holomorphic global coordinates where
It is known that is a non-compact complete Kähler manifold of constant sectional curvature under this metric. From (29), the associated Laplace-Beltrami operator is
Take as a reference point. The associated Poincaré distance function of from 0 on can be easily written as
Hence, we have the minimal positive global Green function of for with a pole at 0:
where denotes the standard Euclidean area of the unit sphere in By definition (see Section 2.1), we have
This gives
which is just the geodesic ball centered at 0 with radius in Hence, the boundary of is the geodesic sphere centered at 0 with radius Whence, we obtain the Green function of for with a pole at 0 satisfying Dirichlet boundary condition:
So, for with
We write
in the (geodesic) polar coordinate form, where is the Riemannian metric on induced by and
is the standard Euclidean metric on It yields from (29) that
where
Let be the Riemannian area element of Using the second formula in (31) and volume form formula, it gives immediately
where is the standard Euclidean area element of
With these preparations, we shall give an estimate for Ricci term Let be the scalar curvature of which is a negative constant. It follows from (32) that
Substituting (30) into the expression of and calculating the integral, it is not hard to deduce that
We proceed to estimate with By
we have Hence, we obtain:
Corollary 5.8.
Let be a nonconstant -valued algebroid function on Equip with the Poincaré metric. Then
holds for all outside a subset of finite Lebesgue measure.
6. Unicity Problems of Algebroid Functions
Assume that is a non-compact complete Kähler manifold.
6.1. Propagation of Algebraic Dependence
Let be the Fubini-Study form on Given a -valued algebroid function on For define
It is not hard to deduce that Since and we obtain
Fix any integer A proper algebraic subset of is said to be decomposible, if there exist positive integers with for some integer and algebraic subsets with such that If is not decomposable, then we say that is indecomposable. Given algebroid functions Put
Definition 6.1.
Let be an analytic subset of We say that nonconstant algebroid functions on are algebraically dependent on if there exists a proper indecomposable algebraic subset of such that If so, are said to be -related on
Let be an effective divisor on and fix any positive integer (allowed to be ). Write for distinct irreducible hypersurfaces of and positive integers Define the support of with order at most by
It is clear that Set
Let be hypersurfaces of satisfying that if or if for Let be distinct values in and be positive integers (allowed to be ).
We introduce the notation as follows.
When is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), denote by
the set of all such that
When has non-positive sectional curvature satisfying (3), denote by
the set of all such that
and (5) is satisfied.
Let be the hyperplane line bundle over It gives a holomorphic line bundle over
where is the natural projection on the -th factor with Let for For , define
with
where is a fixed positive -form.
Lemma 6.2.
We have
Proof.
A direct computation leads to
Thus, we have
Apply Dynkin formula, we deduce that
This completes the proof. ∎
Let be the set of all indecomposable hypersurfaces in such that for some Put
Lemma 6.3.
Assume that and for some Then
Proof.
Lemma 6.4.
Let be a -valued algebroid function in Then
Proof.
By we obtain
in the sense of currents, where and are understood as divisors. Thus, we obtain
By this with Theorem 5.4, we obtain
This proves the lemma. ∎
Set
and
where
Theorem 6.5.
Let be a -valued algebrod function in for Assume that are -related on for some If then are -related on
Proof.
Put
Fix any -valued algebroid function We say that is generic with respect to if for at least one Assume that is generic with respect to Let be the set of all such that on Set
Theorem 6.6.
If then contains exactly one element.
Proof.
Let be a -valued algebroid function in It suffices to show that on Set
Let be the diagonal of Since on then Let be the natural projection on the -th factor with Consider the holomorphic line bundle over where is the hyperplane line bundle over If on then there is a holomorphic section of over such that and Taking then we obtain and On the other hand, implies that
where With the help of Theorem 6.5, we obtain which is a contradiction. This proves the theorem. ∎
6.2. Five-Value Type Theorem
A1 has non-negative Ricci curvature
Assume that is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2).
Theorem 6.7.
Let be transcendental -valued, -valued algebroid functions on respectively. Let be distinct values in Then
Assume for all If then
Assume for all If then
In particular, the conclusions hold for nonconstant if
Proof.
Set and with Then
By Theorem 6.6, we have holds. Similarly, when for we deduce that
which implies holds. ∎
Corollary 6.8 (Five-Value Type Theorem).
Let be nonconstant -valued algebroid functions on If share distinct values ignoring multiplicities in then
B1 has non-positive sectional curvature
Assume that has non-positive sectional curvature satisfying (3).
Theorem 6.9.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that satisfy Then
Assume for all If then
Assume for all If then
Proof.
The argument is similar as in the proof of Theorem 6.7. ∎
Corollary 6.10 (Five-Value Type Theorem).
Let be nonconstant -valued algebroid functions on Assume that satisfy If share distinct values ignoring multiplicities in then
A2 has non-negative Ricci curvature
Assume that is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2).
Theorem 6.11.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies then
Proof.
Assume on the contrary that The condition implies that
in the sense of currents, where and are understood as divisors. So, we are led to
It yields from Lemma 6.3 that
Similarly,
Combining (6.2) with (6.2), it yields from Theorem 5.4 that
It is therefore
But, it contradicts with the assumption for This completes the proof. ∎
Corollary 6.12.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Then
Assume for all If then
Assume for all If then
Proof.
Corollary 6.13 (Five-Value Type Theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively. If share distinct values ignoring multiplicities in then
B2 has non-positive sectional curvature
Assume that has non-positive sectional curvature satisfying (3).
Theorem 6.14.
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Assume that for all If satisfies then
Proof.
The argument is similar as in the proof of Theorem 6.11. ∎
Corollary 6.15.
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Then
Assume for all If then
Assume for all If then
Corollary 6.16 (Five-Value Type Theorem).
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. If share distinct values ignoring multiplicities in then