### v11

# Nevanlinna Theory of Algebroid Functions on Complete Kähler Manifolds

Abstract.
In this paper, we generalize the classical Nevanlinna theory of algebroid functions from to a complete Kähler manifold with either non-negative Ricci curvature or non-positive sectional curvature. As its applications, we establish some Picard type theorems and five-value type theorems for algebroid functions under certain conditions.
Key words and phrases:
Algebroid functions; Nevanlinna theory; second main theorem; Picard type theorem.2010 Mathematics Subject Classification:
32H30; 32H25Contents
- 1 Introduction
- 2 Algebroid Functions on Hermitian Manifolds
- 3 Uniformization of Algebroid Functions without Construction
- 4 Nevanlinna Theory of Meromorphic Functions
- 5 Nevanlinna Theory of Algebroid Functions
- 6 Unicity Problems of Algebroid Functions
1. Introduction
1.1. Motivation
The notion of algebroid functions (see [16, 28]) introduced by H. Poincaré is an extension of meromorphic functions. Recall that a -valued algebroid function on a domain is defined by an irreducible algebraic equation
where are holomorphic functions without common zeros in An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G. Darboux deemed that it is an important class of functions. To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a -sheeted ramified analytic covering of A further fact states that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
The most concerned problem on algebroid functions seems to be the value distribution. In the 1930s after R. Nevanlinna [22] founded two fundamental theorems concerning the value distribution theory of meromorphic functions in 1925, many famous scholars such as G. Rémoundos, H. Selberg, E. Ullrich and G. Valiron, etc. began to study the value distribution theory of algebroid functions. For example, G. Rémoundos [24] proved a Picard’s theorem which claims that a nonconstant -valued algebroid function on omits distinct values at most in Afterwards, G. Valiron [40] showed without details a five-value type theorem in unicity theory, which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to the study of Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc. for algebroid functions. We refer readers to H. Cartan [7], He-Xiao [16], Niino-Ozawa [19], J. Noguchi [20], M. Ozawa [23], M. Ru [25, 26], Sun-Gao [28], J. Suzuki [29], L. Selberg [30, 31], N. Toda [37, 38] and E. Ullrich [39], etc.
Algebroid function theory in one complex variable didn’t mature until the last 30-40 years. Up to now, there seems to be very little literature regarding algebroid functions in several complex variables, except for the limited work by Hu-Yang [15]. Let denote any two non-constant algebroid functions on a non-compact Hermitian or Kähler manifold We are inspired by the following three fundamental problems:
How to realize the uniformization of ?
How many Picard’s exceptional values can have at most?
How many values should share at least to guarantee ?
In the present paper, we answer these problems. What is more important is that we develop the classical Nevanlinna theory of algebroid functions on to complete non-compact Kähler manifolds.
1.2. Main Results
A. Uniformization
Let be a non-compact Hermitian manifold. We generalize the notion of algebroid functions from to (see Definition 2.15). Let be a -valued algebroid function on Given a -leaf function element of When is not a singular branch point of we give a local Pusieux expansion of at i.e., in a local holomorphic coordinate we prove that can be expanded locally into the following convergent Pusieux series (see Theorem 2.14)
at with where are meromorphic functions. By this Pusieux series, we construct a -leaf complex manifold and show that
Theorem A.
Let be the -leaf complex manifold by a -valued algebroid function on Then, can lift to a meromorphic function on via the natural projection such that
B. Nevanlinna theory
In this part and the following Parts C, D, we use to stand for a complete non-compact Kähler manifold. Let be any -valued algebroid function on Refer to the definition in Section 5.3, we have the Nevanlinna’s functions of and on where is the Ricci form of and is a precompact domain containing a fixed point To unify the form, we write
The simple defect of with respect to is defined by
(b1) has non-negative Ricci curvature
Assume that is non-parabolic. Set
where denotes the Riemannian volume of the geodesic ball centered at with radius
Theorem B1 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
When we have Hence, Theorem B1 covers the classical second main theorem for with (see Corollary 5.5).
Corollary B2 (Defect Relation).
Assume the same conditions as in Theorem B1. Then
if one of the following conditions is satisfied
satisfies the volume growth condition
where is defined by
is of non-polynomial type growth, i.e., satisfies the growth condition
Note that as due to
The condition (2) is relaxed. For example, it is satisfied when with or when with or and
Corollary B3 (Picard Type Theorem).
Let be a nonconstant -valued algebroid function on Then, can omit distinct values at most in if one of the conditions in Corollary is satisfied.
(b2) has non-positive sectional curvature
Assume that has curvature satisfying
for constants with that either or where are the sectional curvature and Ricci curvature of respectively. Set
Theorem B4 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
Note that Theorem B4 covers the classical second main theorem for both and the unit ball with Poincaré metric in (see Corollaries 5.7 and 5.8).
Corollary B5 (Defect Relation).
Assume the same conditions as in Theorem B4. If
then
Corollary B6 (Picard Type Theorem).
Let be a nonconstant -valued algebroid function on satisfying Then, can omit distinct values at most in
C. Propagation of algebraic dependence
Propagation problem for algebraic dependence of meromorphic mappings was first studied by L. Smiley [32] and further studied by Y. Aihara [3, 4, 5], S. Drouilhet [9], H. Fujimoto [13, 14], Dulock-Ru [10] and W. Stoll [36], etc.
Assume that is either non-parabolic with non-negative Ricci curvature, or of curvature satisfying (3).
Let be an effective divisor on where are irreducible hypersurfaces of Let be a positive integer (allowed to be ). Set
Let be hypersurfaces of satisfying that if or if for Let be distinct values in and be positive integers (allowed to be ). Denote by the set of all nonconstant algebroid functions on Set
Let
be the set of all such that
and that either (2) is satisfied if is non-parabolic with non-negative Ricci curvature, or (5) is satisfied if has curvature satisfying (3).
Let be the point line bundle over We denote by the set of all indecomposable hypersurfaces of such that for some in which and is the natural projection on the -th factor. Set
with and
Theorem C1.
Let be a -valued algebrod function in for Assume that are -related on for some If then are -related on
Let be a positive integer. Set
Fix any Assume that for at least one Let be the set of all such that on Put
Corollary C2.
If then contains exactly one element.
D. Unicity problems
(d1) has non-negative Ricci curvature
Assume that is non-parabolic with volume growth satisfying (2). Let be a positive integer. Consider the following condition
with
Theorem D1.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies then
Corollary D2 (Five-Value Type Theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively. If share distinct values ignoring multiplicities in then
(d2) has non-positive sectional curvature
Assume that has curvature satisfying (3).
Theorem D3.
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Assume that for all If satisfies then
Corollary D4 (Five-Value Type Theorem).
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. If share distinct values ignoring multiplicities in then
2. Algebroid Functions on Hermitian Manifolds
2.1. Notion of Algebroid Functions
First of all, we would like to extend the notion of algebroid functions to a non-compact Hermitian manifold. Let be an -dimensional non-compact Hermitian manifold. Consider an irreducible algebraic equation in on :
where are holomorphic functions locally defined on such that they well define a meromorphic mapping
Let be the indeterminacy locus of which is either an analytic set of complex codimension not smaller than for or an empty set for in Note that is holomorphic on
Remark 2.1.
The local definition for means that i.e., a collection of holomorphic functions for all such that
where is an open covering of Whence, we see that is well defined on and it hence defines a meromorphic function on with
2.1.1. Meromorphic function elements
Set
It gives that
Definition 2.2.
Let be any point in A pair is called a regular function element, if is a holomorphic function in a neighborhood of and called a meromorphic function element, if is a meromorphic function in a neighborhood of A pair is said to be subordinate to equation if is a local solution of equation in a neighborhood of
Theorem 2.3.
Let be a pair such that
Then, there exists a unique regular function element subordinate to equation satisfying
Proof.
The theorem can be confirmed by applying implicit function theorem. Here, we give an alternative proof in complex analysis. Take a such that given by Remark 2.1. Write as the form
where
By the conditions, we have
The continuity implies that here exist such that
for all with and In further, restrict to then we obtain
According to Rouché’s theorem, we deduce that for all with and have the same number of zeros in the disc In fact, this number of zeros is 1, since in has the unique zero in the disc for all with It implies that uniquely determines a function in a smaller geodesic ball neighborhood of with radius such that Evidently, is continuous on
Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking a local holomorphic coordinate near such that Thanks to Hartog’s theorem, it is sufficient to examine that is holomorphic in each holomorphic direction at For and put
Note that
as That is,
Thus, is holomorphic in the holomorphic direction at This completes the proof. ∎
Definition 2.4.
A point is called a critical point, if either or has a multiple root as a polynomial in We use to denote the set of all critical points, called the critical set. A regular point means a point in where is called the regular set. Moreover, denote by the set of all points in such that has a multiple root as a polynomial in where is called the multiple set and a point in is called a multiple point.
For any the fundamental theorem of algebra asserts that
has exactly distinct complex roots. Hence, Theorem 2.3 implies that
Corollary 2.5.
For any there exist exactly distinct regular function elements subordinate to equation
To see the common roots of and we define the resultant of and by a -order determinant:
where for Apply Sylvester’s elimination method, we can deduce that if and only if or has a multiple root as a polynomial in i.e., Thus, it means that is an analytic set of complex codimension Define the discriminant of by
Assume that splits into linear factors, say
A direct computation leads to
If then has a multiple root as a polynomial in If we set and rewrite (7) as where
Since we see that is a zero of i.e., a pole of In fact, we note from Theorem 2.22 (see Section 2.3) that if and only if To conclude, Corollary 2.5 can be extended to as follows.
Corollary 2.6.
For any there exist exactly distinct meromorphic function elements subordinate to equation
2.1.2. Analytic continuation of meromorphic function elements
We recall that a meromorphic function element on is a pair such that is a meromorphic function defined in a neighborhood of For convenience, we write without any confusion. We usually take which is a geodesic ball neighborhood centered at with radius For two meromorphic function elements and define an equivalent relation: if and only if
Set
We equip with a topology. A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of By the analytic continuation, we may regard that
A neighborhood of is defined by a set (containing ) of all such that is a direct analytic continuation of satisfying where is a neighborhood of in That is,
In particular, a -neighborhood of (satisfying that ), is defined by the set (which contains ) of all such that is a direct analytic continuation of satisfying In other words, it is defined by A subset is called an open set, if either is an empty set or every point is an inner point, i.e., there is a -neighborhood of such that It is not hard to check that it defines a topology of
Theorem 2.7.
is a Hausdorff space.
Proof.
Pick If then we can consider -neighborhoods of and of where Evidently, we have If for then we can take a function element which gives that in a smaller neighborhood of Using the uniqueness theorem of analytic functions, we obtain a contradiction with ∎
Definition 2.8.
Let be two meromorphic function elements on Let be a curve with and We say that is an analytic continuation of along if
for any there is a meromorphic function element on
for any and any there exists such that for all and that is a direct analytic continuation of
Remark 2.9.
The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Moreover, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve, i.e., the analytic continuation is symmetric.
Theorem 2.10 (Analytic continuation).
Let be a curve. Then, a function element can continue analytically to along and attain to a unique function element
Proof.
Set
Evidently, It is sufficient to prove that Otherwise, we may assume that By Corollary 2.6, there exist exactly function elements at denoted by
Take small enough such that with Then, there is a unique integer such that Notice that the domain of definition of is for each one can redefine function element Combined with the definition of it yields immediately that can continue analytically to along i.e., is an analytic continuation of along However, it contradicts with the definition of Hence, we obtain The uniqueness of is immediate due to the uniqueness theorem of analytic functions. ∎
Corollary 2.11.
Let be any two groups of distinct function elements at respectively. Then, can continue analytically to each other along any curve in
Theorem 2.12 (Connectivity).
Any two function elements in can continue analytically to each other along a curve in
Proof.
Let be distinct function elements at every point With the help of Corollary 2.11, it is sufficient to prove that all can be continued analytically to each other. Let be the largest integer such that there are function elements in which continue analytically to each other along a curve. Note that Without loss of generality, one may assume that the function elements are just Then, it can be written as
where are holomorphic functions in a neighborhood of Using Vieta’s theorem, we obtain
It suffices to show that
can continue analytically to the whole
Apply Theorem 2.10, can continue analytically to along a curve for any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., can continue analytically to Hence, holds. If then one can determine another set using the rest function elements However, it leads to that equation (7) is reducible, which is a contradiction. This completes the proof. ∎
Corollary 2.13.
is a path-connected space.
2.1.3. Properties of function elements near a multiple point
Let be a pair such that and
Let us take a small neighborhood of such that on Using the continuity of for there exists a number such that
According to Corollary 2.6, for any with there exist two groups of distinct meromorphic function elements at subordinate to equation (7), respectively. Apply Theorem 2.10, then for a curve with and we are able to extend analytically to along In other words, there is a permutation of such that can continue analytically to along respectively. The symmetry implies that can also continue analytically to along However, we should mention that won’t be, in general, continued analytically back to along respectively. In what follows, we would consider the analytic continuation of a function element along a closed Jordan curve surrounding any multiple point Take a smooth closed Jordan curve around such that which is close to and doesn’t surround or cross any connected components of that do not pass through Without loss of generality, we look at the analytic continuation of along starting from When returns to along it will turn to certain function by which we mean that for a local coordinate chart ( is small enough), jumps to some single-valued component when travels around once along started at due to the change of -dimensional argument of If then one continues to extend analytically along in the same direction of rotation. Repeat this action, we shall see finally that cycles on and on periodically. Let be the smallest period such that can return to itself along all such curves Then, we can obtain We call a branch point with order of if Repeatedly, it can be seen that is divided into finitely many groups. Let be the number of groups. If the -th group has members for then
We just analyze the first group: Taking a suitable there exists a biholomorphic mapping such that
where is the polydisc centered at 0 with polyradius in Setting where is the unit disc centered at 0 in the -plane. For any we have
Set and write
We show that is a single-valued analytic function on Notice that is the branch order of we deduce that is holomorphic on because is so close to 0 that there exist no any branch points of in We only need to examine that is a removed singularity of To do so, it suffices to prove that is bounded near 0 due to Riemann’s analytic continuation. In fact, it follows from (7) and (9) that
on If we obtain Thus, on i.e., is bounded on It means that is holomorphic on and it has the Taylor expansion
at Let be the set of all branch points of in Saying that a branch point of is non-singular, if does not lie in any crossings of components of and singular otherwise. Assume that is not a singular branch point of Then, is smooth when is small enough. Taking a suitable if necessary. Since is bounded on and defines the branch order of at that is independent of the choice of we can conclude from and Riemann’s analytic continuation that there exists a biholomorphic mapping with
such that can be expanded locally into the following Pusieux series
at where are holomorphic functions in Comparing coefficients, we deduce that
According to the above arguments, we show that is a -valued analytic function defined in a neighborhood of We call a -leaf function element. If one cuts along a real hypersurface passing through then can separate into single-valued analytic components, and all these components can turn to each other when they continue analytically along a closed curve around in that is sufficiently close to By the local expansion of we note that is a -multiple, -valued branch point with order of
Assume that is not a singular branch point of We set and define
By the previous arguments, one can also expand locally into a Pusieux series of the form
at If then can be expanded locally at in the form
We take a number with Set Then
It is clear that
Hence, the case turns to the case
To conclude, for each there are distinct multi-leaf function elements subordinate to equation (7), denoted by where is a -valued meromorphic function with and Hence, has branch order at with If then we say that is a branch point with order of Furthermore, if is not a singular branch point of then we deduce that can expand locally into the convergent Pusieux series at That is, we have shown that
Theorem 2.14.
Let be a -leaf function element subordinate to equation If is not a singular branch point of then can be expanded locally into the following convergent Pusieux series
at in a local holomorphic coordinate such that and for all where is the branch set of and are meromorphic functions.
Note that in Theorem 2.14 is a -multiple root of that is, can separate into distinct single-valued components with value at If then separates into distinct single-valued components which take at
According to the analytic continuation and uniqueness theorem, the above arguments show that equation (7) uniquely defines a -valued meromorphic function on outside the branch points of Let stand for the set of all branch points of which is an analytic set in called the branch set of We extend the notion of algebroid functions from to as follows.
Definition 2.15.
A -valued algebroid function on is defined by an irreducible algebraic equation
where are holomorphic functions locally defined on such that they define a meromorphic mapping
In a word, denote by the distinct single-valued components of and given a -leaf function element subordinate to equation (7), with distinct single-valued components Then, there exists a permutation of with for on
In the end, we provide an example to illustrate how to calculate the order of branch points of algebroid functions.
Example. Calculating the branch order of the algebroid function
on at branch points
Since have single-valued components, respectively, we infer that has single-valued components; Since have single-valued components, respectively, we infer that has single-valued components. Thus, is a -valued algebroid function. Let stand for any closed Jordan curve in All the branch sets of are as follows
It is clear that
Order of
Take such that is surrounded by Note that are the only two branch sets of that pass through So, is not allowed to surround or pass through any other branch sets of except for in order to determine the order of For all such we for
Hence, the order of is determined by the orders of branch point of and the order of branch point of It is
Order of
Take such that is surrounded by Note that are the only two branch sets of that pass through Similarly, cannot surround or pass through any other branch sets of except for It implies that for
Hence, the order of is determined by the orders of branch points of It is
Order of
Take such that is surrounded by Since are the only three branch sets of that pass through we have for
Hence, the order of is determined by the orders of branch points of respectively. It is
Order of
Take such that is surrounded by Since are the only three branch sets of that passes through we have for
Hence, the order of is determined by the orders of branch points of It is
2.2. Multi-leaf Complex Manifolds by Algebroid Functions
Let A neighborhood of is defined to be a set of all with such that is a direct analytic continuation of i.e.,
Take a suitable we have a biholomorphic mapping such that where is the unit ball centered at the coordinate origin 0 in Dfine a natural mapping by
Clearly, is a homeomorphism whose inverse is
Set
We prove that is a complex atlas of Let It suffices to examine that the transition mapping
is a biholomorphic mapping. This is automatically a homeomorphism. Note that for each we have in a small neighborhood of i.e., A direct computation gives that
for any Similarly, we have
It shows that is a biholomorphic mapping. On the other hand, each corresponds to distinct function elements That is, we conclude that
Theorem 2.16.
is a non-compact -leaf complex manifold of complex dimension
Set
Corollary 2.17.
is a non-compact -leaf complex manifold of complex dimension which is a submanifold of
In what follows, we would construct a -leaf complex manifold determined by all function elements subordinate to equation (7). Let be any two function elements subordinate to equation (7). We define an equivalent relation as follows: if and only if
for some two single-valued components of respectively, where is the branch set of with Set
Definition 2.18.
A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branch set of If so, one may regard that
Let A neighborhood of is a set which contains the following elements:
all with satisfying that is a direct analytic continuation of
all with Here and before, is an arbitrary neighborhood of
In particular, a -neighborhood of with is defined by the set that contains the following elements:
all with satisfying that is a direct analytic continuation of
all with
A subset is called an open set, if is an empty set or every point is an inner point, i.e., there exists a -neighborhood of such that It gives a topology of which contains the topology of If is a -leaf function element, then has leaves at We call a -leaf neighborhood of
Theorem 2.19.
is a connected Hausdorff space.
Proof.
The Hausdorff property of is showed using the similar arguments as in the proof of Theorem 2.7. We assert that is connected. Otherwise, one may assume that where are two disjoint open subsets of Set It is evident that Since is an analytic set of complex codimension we see that are two disjoint open subsets of This implies that is not connected, which is a contradiction. ∎
Next, we will equip with a complex structure such that is a complex manifold. Taking a -leaf function element Assume that is not a singular branch point of Let be chosen suitably. By means of Theorem 2.14, there exists a biholomorphic mapping such that can be expanded into the Pusieux series
on with and for all
Let be a -leaf neighborhood of containing the following elements:
all with satisfying that is a direct analytic continuation of
all with
Let be distinct single-valued components of Define a mapping by
for , where We note that is a homeomorphism and its inverse is
where is an integer such that
Let stand for the set of all function elements such that is a singular branch point of Moreover, we denote by the set of all the corresponding singular branch points of which is obvious an analytic set of complex codimension not smaller than Set
We now show that is a complex atlas of Assume that It is sufficient to verify that the transition mapping
is a biholomorphic mapping. It is automatically a homeomorphism. Through a simple analysis, we only need to deal with the situation in which is a -leaf meromorphic function element and is a regular function element. For any there exists a single-valued component of satisfying that in some neighborhood of i.e., It is clear that since all function elements in are meromorphic. A direct computation leads to
where is an integer such that
Thus, is a holomorphic mapping. On the other hand, its inverse is
where is the -th single-valued component of Evidently, is also a holomorphic mapping. Hence, is a non-compact -leaf complex manifold of complex dimension
Finally, we shall treat the singular branch set Let be the natural projection which maps a function element to Note that is a -sheeted ramified analytic covering such that
Since is an analytic set of complex codimension not smaller than the complex structure of can be viewed as a natural extension of the complex structure of This implies that one can naturally extend the complex structure of to via Hence, we conclude that
Theorem 2.20.
is a -sheeted ramified analytic covering of We call the -leaf complex manifold by
Corollary 2.21 (Uniformization).
An algebroid function on can lift to a meromorphic function on via the natural projection by mapping to such that
2.3. Zero and Pole Divisors of Algebroid Functions
Let be a -valued algebroid function on defined by equation Let be single-valued components of
Theorem 2.22.
We have
Proof.
Let be a zero of Since in the sense of divisors, we may assume that It follows that
where denotes the number of -valued points of a function at Without loss of generality, we assume that are all components that take at According to Vieta’s theorem, we obtain
It implies that since reaches the maximal pole order at It yields from (11) that
Thus, we conclude that ∎
Similarly, we can show by setting that
Theorem 2.23.
We have
3. Uniformization of Algebroid Functions without Construction
In this section, we provide another approach to realize the uniformization of algebroid functions without constructing finite ramified analytic coverings. Our main idea originates from W. Stoll [35].
3.1. Algebroid Reduction
Let be a non-compact complex manifold of complex dimension Let be a -sheeted ramified analytic covering, which means a proper, surjective holomorphic mapping with a ramification set i.e.,
is a proper, surjective, local biholomorphic mapping of sheet number and thus a -sheeted analytic covering. Here, is the set of those points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex codimension By R. Remmert [27] (see Andreotti-Stoll [2] also), the exceptional set
is an analytic set satisfying that and Moreover, and for Let The mapping degree of at is defined in such manner: if is a neighborhood of such that then
which is independent of the choice of Note that and if and only if By W. Stoll [35], for any we have
Let be a complex vector space of complex dimension with a basis Let be the projection of where is the natural projection. Let be the -fold symmetric tensor product of i.e.,
where
is called the symmetric tensor product of and stands for the -order symmetric group. Note that is a subspace of -fold tensor product space with Again, put
where for any Naturally, is embedded into by the mapping Thus, we obtain
Definition 3.1.
Let be a nonempty open subset of Let be a continuous mapping. Then, the algebroid reduction of via is a mapping defined by
Theorem 3.2 (Stoll, [35]).
is a continuous mapping. If is holomorphic, then is also holomorphic and extends to a holomorphic mapping
Definition 3.3.
Let be a meromorphic mapping with indeterminacy locus Then, the algebroid reduction of via is a meromorphic mapping with indeterminacy locus defined by
Theorem 3.4 (Stoll, [35]).
Let be a meromorphic mapping with indeterminacy locus Let be the algebroid reduction of via Fix any Assume that there exist a nonempty neighborhood of such that and a reduced representation If is a representation of on then the algebroid reduction of via is a representation of If is reduced, then is also reduced.
According to Theorem 3.4, we obtain the following commutative diagram:
Theorem 3.5 (Stoll, [35]).
Let be a differentiably non-degenerate and irreducible meromorphic mapping. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via
3.2. Uniformization
In what follows, we consider the case when For convenience, we write and for short. Take a basis of For any there is a unique representation
for certain It induces a natural isomorphism defined by
which follows that If then (13) defines an algebraic equation
Let be the maximal integer in such that Then, equation (14) reduces to Let be the roots of the equation. Applying Vieta’s theorem, it is immediate that
So, is decomposable, i.e.,
By this with (12), we have Hence, we obtain
Lemma 3.6.
Let be a differentiably non-degenerate and irreducible meromorphic mapping. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function which separates the fibers of such that is the algebroid reduction of via
Theorem 3.7 (Uniformization).
Let be an algebroid function. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function such that
Proof.
Let be defined by equation (7). It gives a differentiably non-degenerate, irreducible meromorphic mapping
Set with a basis By (15), we can write
in the sense of isomorphisms. In further, the Vieta’s theorem leads to
in the sense of isomorphisms. On the other hand, Lemma 3.6 says that there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via Hence, we have
with
Let be the -valued points of Then
Combining with then we prove the theorem. ∎
4. Nevanlinna Theory of Meromorphic Functions
Let be a complete non-compact Kähler manifold of with Laplace-Beltrami operator The Kähler form of is defined by
in a local holomorphic coordinate Moreover, the Ricci form of is defined by
4.1. First Main Theorem
Let be any precompact domain with smooth boundary in Fix a point and denote by the Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
where is the Dirac’s delta function on with a pole at The harmonic measure on with respect to is defined by
where is the Riemannian volume element of and is the inward normal derivative on
Next, we introduce Nevanlinna’s functions. Let be a meromorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and with
The characteristic function of on with respect to is defined by
Let The proximity function of on with respect to is defined by
where denotes the spherical distance on defined by
with
The counting function of on with respect to is defined by
We need the following Dynkin formula (see, e.g., [1, 11, 12]).
Lemma 4.1 (Dynkin Formula).
Let be a -class function on outside a polar set of singularities at most. Assume that Then
Theorem 4.2 (First Main Theorem).
Assume that Then
Proof.
Applying Dynkin formula to we get
Note that the first term on the left hand side equals and the right hand side equals
This proves the theorem. ∎
4.2. Second Main Theorem for Non-negative Ricci Curvature
Assume that is non-parabolic. Fix a reference point and denote by the geodesic ball centered at with radius in Let denote the Riemannian volume of The non-parabolicity of means that
Thus, we have the unique minimal positive global Green function of for We can write as
in which is the heat kernel of on Let be the Riemannian distance function of from By Li-Yau’s estimate [18], there exist constants such that
holds for all
4.2.1. Construction of and Nevanlinna’s functions
For define
Since
we see that is a precompact domain containing with and Moreover, the family can exhaust i.e., for any sequence with we have
So, the boundary of can be formulated as
By Sard’s theorem, is a submanifold of for almost all
Set
Note that defines the Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
Let denote the harmonic measure on with respect to defined by
where is the inward normal derivative on is the Riemannian area element of
Let be a meromorphic function. Take in Section 4.1, we have the Nevanlinna’s functions of :
4.2.2. Gradient estimate of Green functions
Lemma 4.3.
For we have
holds for all where is given by
Proof.
Invoking the definition of Green function for it is immediate to deduce that for
Since for we obtain
Let stand for the gradient operator on a Riemannian manifold. Cheng-Yau [8] proved the following theorem.
Lemma 4.4.
Let be a complete Riemannian manifold. Let be a geodesic ball centered at with radius Then, there exists a constant depending only on the dimension of such that
holds for any non-negative harmonic function on where is the Riemannian distance between and and is the lower bound of Ricci curvature of
Theorem 4.5.
There exists a constant such that
Proof.
Applying
we obtain an upper bound estimate of as follows.
Corollary 4.6.
There exists a constant such that
where is the Riemannian area element of
4.2.3. Calculus Lemma and Logarithmic Derivative Lemma
Lemma 4.7 (Borel’s Growth Lemma).
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
The conclusion is clearly true for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit exists or equals If then Set
Note that exists for almost all Then, we have
This completes the proof. ∎
Set
Theorem 4.8 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then there exists a constant such that for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Invoking Lemma 4.3, we have
Set
A simple computation leads to
In further, we have
Applying Borel’s growth lemma to the left hand side of this equality twice: one is to and another is to we deduce that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.6 implies that there exists a constant such that
Set Combining the above, we have
holds for all outside ∎
In the following, we establish a logarithmic derivative lemma. Let be a meromorphic function on The norm of the gradient of is defined by
in a local holomorphic coordinate where is the inverse of Define the Nevanlinna’s characteristic function of by
where
It is not difficult to show that
On one puts a singular metric
with
Lemma 4.9.
We have
Proof.
Note that
It follows from Fubini’s theorem that
By this with the First Main Theorem, we conclude that
Lemma 4.10.
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Define
Theorem 4.11 (Logarithmic Derivative Lemma).
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Note that
By Lemma 4.18, we have for any there exists a subset of finite Lebesgue measure such that
holds for all outside ∎
4.2.4. Second Main Theorem
The characteristic function of is defined by
Theorem 4.12 (Second Main Theorem).
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Define a non-negative function on by
where
is a singular volume form on By we obtain
in the sense of currents, in which is the zero divisor of and is the ramification divisor of Using Dynkin formula, we have
On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on and on as well as
for On each write
where is a positive smooth function on Given a partition of the unity subordinate to Put
Again, set Then on we have
For any we can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
It yields that
Put Then, we obtain
on It is therefore
on Since is bounded on then it yields that
By this with Dynkin formula and Theorem 4.11, for any there exists a subset of finite Lebesgue measure such that
holds for all outside in which we used the fact that Combining (4.2.4) with (20), we have the theorem proved. ∎
4.3. Second Main Theorem for Non-positive Sectional Curvature
Assume that has curvature satisfying (3). According to the definition of curvature, it is not difficult to see that In particular, we have when has constant sectional curvature Fix a reference point We may assume without loss of generality that is a pole of for a technical reason if necessary, since, otherwise, one can consider the universal covering If is equipped with pullback metric induced from via then preserves the same curvature as Note that is a Cartan-Hadamard manifold. In the following, we use to denote the Riemannian distance function of from on
4.3.1. Construction of and Nevanlinna’s functions
A. Setups for
By the curvature conditions, there exists a unique minimal positive global Green function of with a pole at for Consider the following initial value problems on :
and
Evidently, the two ODEs have the unique solutions
Using the estimates of we have (see [17, 34])
where
Here, is the standard Euclidean area of For define
Since as and one infers that is a precompact domain containing in which exhausts Thus, the boundary of can be formulated as
Set
which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. Denote by the harmonic measure on with respect to i.e.,
where is the inward normal derivative on is the Riemannian area element of In particular, if In this case, we have
which is just the geodesic ball centered at with radius in Then
Let be the geodesic ball centered at with radius in We give the relation between and as follows.
Set
Theorem 4.13.
For we have and
Hence, there exists a constant independent of such that
Proof.
B. Setups for
We shall extend the settings for to the case that For the function does not make sense. However, since
we can extend continuously to defined by (4).
It is noted that is parabolic if i.e., there exist no positive global Green functions for In other words, for Hence, one cannot use to define the domain in this case. However, we note (22) for it is thus natural to define
for Let then it yields from (23) that
To conclude, has the unified form
for In further, we have
for
4.3.2. Calculus Lemma and Logarithmic Derivative Lemma
Lemma 4.14.
For we have for
Proof.
When is the geodesic ball centered at with radius The conclusion holds immediately due to (4.3.1). In the following, we consider the other case. Write
Since on we deduce that
Hence, for
This completes the proof. ∎
Theorem 4.15.
There exists a constant such that
holds for all
Proof.
When is a geodesic ball centered at with radius Hence, it yields from (4) and (4.3.1) that
This implies the conclusion holds. For by Lemma 4.4 with (3), (4) and (21) (see Remark 5 in [34] also)
for some large constant which depends only on the dimension By (21) again
which implies that for all Thus, we conclude that
where ∎
Corollary 4.16.
There exists a constant such that
holds for all where is the Riemannian area element of
Set
Theorem 4.17 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then for any there exists a set of finite Lebesgue measure such that
holds for all outside where is defined by
Proof.
Lemma 4.14 yields that
Differentiating to get
In further, we have
Applying Borel’s growth lemma to the left hand side on this equality twice, we can deduce that there exists a set of finite Lebesgue measure such that
holds for all outside Meanwhile, Corollary 4.16 means that there exists a constant such that
Substituting (28) into (4.3.2), we conclude that
for all outside with ∎
Lemma 4.18.
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
Proof.
Theorem 4.19 (Logarithmic Derivative Lemma).
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
Proof.
4.3.3. Second Main Theorem
Theorem 4.20 (Second Main Theorem).
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
Proof.
Applying the similar arguments as in the proof of Theorem 4.12 and combining Theorem 4.17 with Theorem 4.19, we conclude that for any there exists a subset of finite Lebesgue measure such that
holds for all outside
In the following, we estimate When we have
This implies that the theorem holds. We consider the other case that Since as we deduce that
and
Thus, we conclude that
This completes the proof. ∎
5. Nevanlinna Theory of Algebroid Functions
Let be a complete non-compact Kähler manifold of with Laplace-Beltrami operator associated to the Kähler metric We utilize the same notations such as etc. defined in Section 4.
5.1. First Main Theorem
Let denote a precompact domain with smooth boundary in Fix a point As defined in Section 4.1, we have the Green function for and the harmonic measure on
In what follows, we introduce Nevanlinna’s functions of algbroid functions. Let be a -valued algebroid function with branch set on which is defined by an irreducible algebraic equation
such that defines a meromorphic mapping with indeterminacy locus In the following, we assume that
The characteristic function of on is defined by
where
are called the proximity function, counting function of on respectively. For define
where
In order to unify the form, we write
Theorem 5.1 (First Main Theorem).
Assume that Then
Proof.
According to Vieta’s theorem, we have
Since is polar, it yields from Dynkin formula that
By the definition, the left hand side equals
On the other hand, Theorem 2.22 and Theorem 2.23 imply that the right hand side equals
Then, we arrive at
Setting to get
where
In particular, we have
Similarly, we obtain
which yields that
Since
we deduce that
Put together the above, we conclude that
5.2. Estimate of Branch Divisors
Let be a -valued algebroid function on We use the same notations as defined in Section 5.1. Our purpose is to obtain an estimate of the branch divisor of For each we use to stand for the all -valued, -valued components of at respectively. Then, we have and has branch order at with and has branch order at The branch divisor of associated to is defined by a divisor on such that
The counting function of branch divisor of for is defined by
where
Recall that the discriminant of is given by
Lemma 5.2.
We have
Proof.
For any we use to denote the all distinct function elements of at with leaves respectively. It is noted that From the assumption, we deduce that Since the set of all singular branch points of has complex codimension not smaller than it is sufficient to deal with the case where is a non-singular branch point of By Theorem 2.14, we may write for
in a local holomorphic coordinate near such that has a local holomorphic coordinate If then is not a pole of Through a simple observation of we deduce from (29) immediately that there exist terms that contain the factor i.e., terms that contain the factor in For the worst case where for all we deduce that the factor appears a total of at least times in Hence, we get
If then is a pole of For the worst case where for all it yields from Theorem 2.22 that
Combining the above two cases, we prove the lemma. ∎
Set
Theorem 5.3 (Estimate of branch divisors).
We have
5.3. Second Main Theorem
Let be the -leaf complex manifold by a -valued algebroid function on which is of complex dimension by Theorem 2.20. Let be the natural projection. We equip with the pullback metric induced from via Note that is a positive semi-definite Kähler metric on i.e., is a Kähler metric on and degenerate on Since is a polar set with Dynkin formula still works on for the following reasons. Set Picking a sequence of neighborhoods of such that
Let be a partition of the unity subordinate to the open covering of Then, and Fix a Kähler metric on with Kähler form Set Since is positive definite on we see that is also positive definite on Moreover, since we conclude that is a Kähler metric on with on Let denote the Riemannian measures of respectively. By and as we can deduce that weakly as Let stand for the Laplace-Beltrami operators of on respectively. According to and weakly as one has weakly as Taking a bounded domain with smooth boundary in Fixing with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, here is well-defined on Also, we denote by the harmonic measures of on with respect to respectively, here is well-defined on Since is a polar set, the previous arguments can show that in the sense of distributions and weakly as By Dynkin formula
for a -class function with on outside a polar set of singularities at most. Since is polar, we conclude by letting that
This means that Dynkin formula works on under the pullback metric By Corollary 2.21, one can lift to a meromorphic function such that In further, it is not difficult to examine that Theorem 4.2, Theorem 4.12 and Theorem 4.20 can apply to on each leaf of under the corresponding conditions.
5.3.1. has non-negative Ricci curvature
Assume that is non-parabolic. With the pullback metric one notes that is non-parabolic, with non-negative Ricci curvature outside Let be the Kähler form of Set where is the bounded domain defined in Section 4.2.1. We recall the definitions of Green function for and harmonic measure on in Section 4.2.1. Then, the characteristic function of is defined by
where
Indeed, define for
where
In order to unify the form, we write
The simple counting function of with respect to is defined by
Assume that Let be the connected components of centered at respectively, with that for each Let denote the Laplace-Beltrami operator of which is well-defined on Define
In a sense, we can regard outside that is the Green function of for with a pole at satisfying Dirichlet boundary condition, and that is the harmonic measure on with respect to Set
where
Define for
and
If is not constant, then it follows from Theorem 4.12 that for distinct values there is a subset of finite Lebesgue measure such that
holds for and all outside
Theorem 5.4 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Proof.
Note that is a -sheeted analytic covering, and two distinct connected components of meet only at the points in Moreover, any point in is possibly a branch point of or equivalently, any point in is possibly a ramification point of Hence, it follows from that
On the other hand, we have as well as
Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset of finite Lebesgue measure such that
holds for all outside By Theorem 5.3, we prove the theorem. ∎
When we have By it is trivial to deduce that Hence, we obtain:
Corollary 5.5.
Let be a nonconstant -valued algebroid function on with Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
5.3.2. has non-positive sectional curvature
Assume that has curvature satisfying (3). Note that the curvature of also satisfies (3) outside under the pullback metric Similarly, we obtain notations and by setting Employing the similar arguments as in the proof of Theorem 5.4, we can conclude from Theorem 4.20 that
Theorem 5.6 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
When we deduce that and Thus, we conclude that
Corollary 5.7.
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside
Let be the unit ball centered at the origin 0 in Consider equipped with the Poincaré metric
in a global holomorphic coordinate where
It is known that is a complete non-compact Kähler manifold of constant sectional curvature under this metric. Using (31), the associated Laplace-Beltrami operator is
Take as a reference point. The associated Poincaré distance function of from 0 on can be easily written as
Hence, we obtain the minimal positive global Green function of for with a pole at 0:
where stands for the standard Euclidean area of the unit sphere in By the definition (see Section 2.1), we have
This gives
which is just the geodesic ball centered at 0 with radius in Hence, the boundary of is the geodesic sphere centered at 0 with radius Whence, we obtain the Green function of for with a pole at 0 satisfying Dirichlet boundary condition:
Thus, for with
We write
in the (geodesic) polar coordinate form, where is the Riemannian metric on induced by and
is the standard Euclidean metric on It yields from (31) that
with Let be the Riemannian area element of By the second formula in (33), it gives immediately
where is the standard Euclidean area element of
With these preparations, we shall give an upper estimate of Let be the scalar curvature of which is a negative constant. It yields from (34) that
Substituting (32) into the expression of and computing the integral, it is not hard to deduce that We proceed to estimate with By
we have Hence, we obtain:
Corollary 5.8.
Let be a nonconstant -valued algebroid function on Equip with the Poincaré metric. Then
holds for all outside a subset of finite Lebesgue measure.
6. Unicity Problems of Algebroid Functions
Assume that is a complete non-compact Kähler manifold.
6.1. Propagation of Algebraic Dependence
Let be the Fubini-Study form on Given a -valued algebroid function on For define
It is not hard to deduce that Since and we obtain
Fix any integer A proper algebraic subset of is said to be decomposible, if there exist positive integers with for some integer and algebraic subsets with such that If is not decomposable, then we say that is indecomposable. Given algebroid functions Set
Definition 6.1.
Let be an analytic subset of We say that nonconstant algebroid functions on are algebraically dependent on if there exists a proper indecomposable algebraic subset of such that If so, are said to be -related on
We let be an effective divisor on and fix a positive integer (allowed to be ). Write for distinct irreducible hypersurfaces of and positive integers Define the support of with order at most by
Note that Set
Let be hypersurfaces of satisfying that if or if for Let be distinct values in and let be positive integers (allowed to be ).
We introduce the notation as follows:
If is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), then we denote by
the set of all such that
Let be the point line bundle over It gives a holomorphic line bundle over :
where is the natural projection on the -th factor with Let with For , define
with
where is a positive -form.
Lemma 6.2.
We have
Proof.
A direct computation leads to
Thus, we have
Applying Dynkin formula, we conclude that This completes the proof. ∎
We denote by the set of all indecomposable hypersurfaces of such that for some Set
Lemma 6.3.
Assume that and for some Then
Proof.
Lemma 6.4.
Let be a -valued algebroid function in Then
Proof.
Since we obtain
in the sense of currents, where and are understood as divisors. Thus, we are led to
By this with Theorem 5.4, we obtain
This proves the lemma. ∎
Set
where
Theorem 6.5.
Let be a -valued algebrod function in for Assume that are -related on for some If then are -related on
Proof.
Set
Fix any -valued algebroid function We say that is generic with respect to if for at least one Assume that is generic with respect to Let be the set of all such that on Set
Theorem 6.6.
If then contains exactly one element.
Proof.
Let be a -valued algebroid function in It suffices to show that on Set Let denote the diagonal of Since on we obtain Let be the natural projection on the -th factor with We consider the holomorphic line bundle over where is the point line bundle over If on then there exists a holomorphic section of over such that and Take then one can have and On the other hand, implies that
where With the help of Theorem 6.5, we obtain which is a contradiction. This proves the theorem. ∎
6.2. Five-Value Type Theorem
A1 has non-negative Ricci curvature
Assume that is non-parabolic with volume growth satisfying (2).
Theorem 6.7.
Let be transcendental -valued, -valued algebroid functions on respectively. Let be distinct values in Then
Assume for all If then
Assume for all If then
Proof.
Set and with Then
By Theorem 6.6, we have holds. Similarly, when for we deduce that
which implies that holds. ∎
Corollary 6.8 (Five-Value Type Theorem).
Let be nonconstant -valued algebroid functions on If share distinct values ignoring multiplicities in then
B1 has non-positive sectional curvature
Assume that has curvature satisfying (3).
Theorem 6.9.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that satisfy Then
Assume for all If then
Assume for all If then
Proof.
The argument is similar as in the proof of Theorem 6.7. ∎
Corollary 6.10 (Five-Value Type Theorem).
Let be nonconstant -valued algebroid functions on Assume that satisfy If share distinct values ignoring multiplicities in then
A2 has non-negative Ricci curvature
Assume that is non-parabolic with volume growth satisfying (2).
Theorem 6.11.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies then
Proof.
Assume on the contrary that The condition implies that
in the sense of currents, where and are understood as divisors. Thus, it yields from that
In further, it follows from Lemma 6.3 that
Similarly, we have
Combining the above and using Theorem 5.4, we are led to
It is therefore
which contradicts with the condition (6) for This completes the proof. ∎
Corollary 6.12.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Then
Assume for all If then
Assume for all If then
Proof.
It is not hard to show that holds with under the condition for Employing Theorem 6.11, we have holds. Similarly, we have holds with Hence, we prove ∎
Corollary 6.13 (Five-Value Type Theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively. If share distinct values ignoring multiplicities in then
B2 has non-positive sectional curvature
Assume that has curvature satisfying (3).
Theorem 6.14.
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Assume that for all If satisfies then
Proof.
The argument is similar as in the proof of Theorem 6.11. ∎
Corollary 6.15.
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Then
Assume for all If then
Assume for all If then
Corollary 6.16 (Five-Value Type Theorem).
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. If share distinct values ignoring multiplicities in then